### v1

# Professionally translated parallel data for 115 under-represented languages

Abstract
We open-source Smol (Set of Maximal Overall Leverage)111huggingface.co/datasets/google/smol, a suite of training data to unlock translation for low-resource languages (LRLs). Smol has been translated into 115 under-resourced languages, including many for which there exist no previous public resources, for a total of 6.1M translated tokens. Smol comprises two sub-datasets, each carefully chosen for maximum impact given its size: SmolSent, a set of sentences chosen for broad unique token coverage, and SmolDoc, a document-level source focusing on a broad topic coverage. They join the already released Gatitos for a trifecta of paragraph, sentence, and token-level content. We demonstrate that using Smol to prompt or fine-tune Large Language Models yields robust ChrF improvements. In addition to translation, we provide factuality ratings and rationales for all documents in SmolDoc, yielding the first factuality datasets for most of these languages.
SMOL:
Professionally translated parallel data
for 115 under-represented languages
Isaac Caswell∗ Elizabeth Nielsen∗ Jiaming Luo Colin Cherry Geza Kovacs Hadar Shemtov Partha Talukdar Dinesh Tewari Baba Mamadi Diane Koulako Moussa Doumbouya Djibrila Diane Solo Farabado Cissé Google {Research, Deepmind} Stanford University NKo USA INC ∗{icaswell, eknielsen}@google.com
1 Introduction
Most of the world’s 7000 or so languages have few or no resources available for machine translation. However, for an expensive task like professional translation, it is not clear how to make the best use of a limited budget to generate the most effective data. As shown by the Gatitos dataset (Jones et al., 2023), training data consisting of individually translated tokens without context provides large benefits to translation quality for low-resource languages at the lowest cost. However, gains quickly saturate, as single tokens are not very expressive. Sentence-level data is better for a model once token-level data saturates, but it has much more inherent redundancy; and document-level data is even more effective…and more redundant. A resource like Gatitos is a useful first step for the languages it covers, but without sentence- or document-level resources, there is a limit to how effective translation models can be.
In this work, we release the Smol dataset, which provides professionally translated sentence- and document-level data for 115 LRLs, supplementing the token-level translations from Gatitos. Smol contains two sub-datasets:
SmolSent: 863 English sentences covering 5.5k of the most common English tokens,222In this paper, ‘token’ refers to logical/linguistic words, not subword tokens from a model’s vocabulary. professionally translated into 81 languages.
SmolDoc 584 English documents covering a wide range of topics, domains, and tokens, generated by a LLM and professionally translated into 100 languages.
We demonstrate the utility of these data for finetuning and prompting LLMs for translation, and also provide factuality annotations for all documents.
2 Related work
There are several training datasets that comprise parallel data for Low-Resource Languages (LRLs), which we operationally define as any language beyond the first 100 or so supported by most traditional crawls and MT providers. The NLLB-Seed dataset is a sentence-level training set consisting of 6k sentences selected from English Wikipedia and professionally translated into 39 LRLs Team et al. (2022). The Gatitos dataset Jones et al. (2023) represents another approach, consisting of a 4000-entry lexicon, translated into 170 LRLs. There are also several professionally-translated evaluation sets, namely Flores-101 and Flores-200 Goyal et al. (2022); Team et al. (2022), and NTreX (Federmann et al., 2022a).
While highly multilingual, professionally translated training data is rare, there is a growing number of bottom-up community datasources organized through research collectives like Masakhane ( et al., 2020), Turkish Interlingua (Mirzakhalov et al., 2021a, b), and GhanaNLP (Azunre et al., 2021a); and conferences and workshops like AfricaNLP, AmericasNLP (Mager et al., 2021) and ArabNLP. These datasets are usually generated by researchers fluent in the languages, and are therefore especially high quality. In addition to providing datasets, such efforts frequently also provide models and baselines, or even public interfaces, like the Khaya Translator Web App333https://ghananlp.org/project/translator-webapp/ by GhanaNLP for West African languages, and the lesan.ai444https://lesan.ai/translate translation website for Ethiopian languages.
Participation is especially strong from the African continent, including corpora and models for pan East-African languages (Babirye et al., 2022), languages from the Horn of Africa (Hadgu et al., 2022), Ethiopian languages (Teferra Abate et al., 2018; Gezmu et al., 2021), Ugandan languages (Akera et al., 2022), Emakhuwa (Ali et al., 2021), South-African languages (Eiselen and Puttkammer, 2014), Setswana and Sepedi (Marivate et al., 2020), Yorùbá (Adelani et al., 2021b, a), Oshiwambo (Nekoto et al., 2022), Igbo (Ezeani et al., 2020), Zulu (Mabuya et al., 2021), Twi (Azunre et al., 2021b), Gbe (Hacheme, 2021), Bambara (Tapo et al., 2021), and Fon (Emezue and Dossou, 2020). Outside of Africa, corpora have been created for languages of the Americas, including for four indigenous languages of Peru in Bustamante et al. (2020), the numerous results on the largely South- and Central American languages from the first AmericasNLP conference (Mager et al., 2021), and the Inuktitut language of Canada (Joanis et al., 2020). Datasets for lower-resourced languages of India have also sprung up, including the 13-language PMIndia (Haddow and Kirefu, 2020), and datasets focused on languages of the Northeast like Mizo (Thihlum et al., 2020), Khasi (Laskar et al., 2021) and Assamese (Laskar et al., 2020). Finally, a variety of such datasets and models are available for public use on HuggingFace555https://huggingface.co/datasets?multilinguality=multilinguality:translation&task_categories=task_categories:translation&sort=downloads or Zenodo666https://zenodo.org/communities/africanlp/.
3 Methods for getting text to translate
Translation requires significant investment, so great care needs be put into choosing good/effective sentences to translate. This problem is especially tricky because one can’t iterate on the method: once text is translated, the budget is exhausted and one can’t simply select a new source text. It is possible to run ablations from high resource languages, as described in the next section, but it is questionable how much such approaches can actually generalize to true Low-Resource Languages.
For both SmolDoc and SmolSent (as with Gatitos), selection/generation of source text is done in English. This has obvious biases, including the danger of translationese output, a scarsity of translators fluent in English, and a lack of culturally specific domains, but also has large advantages—most notably better scaling for quality control, and the ability to do Researcher in the Loop.
3.1 SmolSent: Token Set Cover
Our basic motivation for creating SmolSent was to help models overcome vocabulary issues, which are common for the lowest-resource languages (Bapna et al., 2022). the alligator problem Nielsen et al. (2025). Therefore, we frame this as a set-cover problem, and pick the smallest set of sentences that cover the largest set of target tokens. The tokens we chose to cover (the target set) were the English side of Gatitos, as well as the most common 2,500 tokens from an English web crawl. Set cover is NP-hard, so we approximate it with a greedy algorithm that iteratively picks the sentence with the highest coverage percent, defined as the percentage of its tokens that are in the target set.
3.1.1 Preliminary work on Token Set-Cover
Method | ChrF |
Random | 30.5 |
Token set-cover | 31.7 |
N-gram DWD | 30.0 |
Embedding DWD | 27.5 |
To evaluate the token set-cover approach, we started by selecting data from existing web-scraped parallel data. We pretrain a multilingual NMT model on parallel data from 294 language pairs from Madlad-400 Kudugunta et al. (2023), with nine languages held out to simulate LRLs. We fine-tune this model on sets of existing parallel data in each of the held-out languages, and evaluate on Flores-200. Details on the experimental set-up can be seen in Appendix A.1.
In addition to Greedy Token Set-Cover, we explore two methods that balance data diversity and data quality. First, we implement Ambati et al. (2011)’s ‘density-weighted diversity’ (DWD) metric, which is an n-gram based metric for diversity and quality. Second, we implement an embedding-based version of DWD, which takes the weighted harmonic mean of perplexity under the Palm 2 model Anil et al. (2023) (proxy for quality), and embedding distance on mBERT sentence embeddings (proxy for diversity). We apply both methods to the English side of the parallel data only, to simulate the case where we don’t yet have LRL translations. As a baseline, we randomly select sentences.
Table 1 shows results after finetuning. Greedy token set-cover performs the best, with diversity-based metrics actively hurting performance.
3.1.2 Researcher in the Loop (RITL)
Despite its success in the ablation, Greedy Token Set Cover had several problems when we scaled it to select from among all the English sentences of CommonCrawl. Firstly, it is maximized by honeypots, or nonsense strings dense in content words (Appendix Table 4); and secondly, it biases towards short sentences, causing length distribution artifacts
These are tricky problems to solve in the general case. However, a dataset like Smol is small enough to manually inspect. Therefore we develop Researcher in the Loop Greedy Set-Cover (Algorithm A.2), where the domain expert (the researcher) can look at and edit each individual sentence.777This work was conducted before the advent of LLMs. Today, this could be simplified using LLMs as autoraters. The result of this process is SmolSent, a set which uses 863 sentences to cover 5519 unique tokens. Qualitatively, SmolSent consists of complex sentences with wide vocabulary coverage; quantitative metrics are explored in Appendix A.3.
3.2 SmolDoc: LLMs with prompt mesh
SmolDoc follows a different and complementary approach. Whereas SmolSent consists of a small set of sentences that are selected from natural text, are complex, and cover many tokens; SmolDoc instead consists of documents that are generated and are simpler, but cover many topics. It should be noted that the token-coverage approach described above failed resoundingly for longer documents, as the power of the honeypots was magnified.
To generate SmolDoc, we used a collection of templates to create a few thousand diverse prompts with a wide range of topics, domains, words, tenses, grammatical cases, and registers (e.g. formal, informal). Appendix B.2 gives details and examples.
Corpus Diversity Ranking for SmolDoc
Document-by-document evaluation as described above does not help one understand corpus diversity — for example, if an almost identical document appears twice, only one of them should be included. Therefore, we rank all candidates by how much new information they add to the corpus, by iteratively finding the document contributing the least new information and removing it, yielding a ranking of all documents. Our criterion for “new information” was the average 9-gram IDF score of a document, minus the fourth moment Bread score (Caswell et al., 2023), to down-weight internally repetitive documents.
Language Tiers for SmolDoc
We wanted to translate text data for languages with more speakers. We break the languages into the five different groups/tiers seen in Appendix Table 6, and for each tier translate the top N documents as ranked by corpus diversity. The smallest tier has 66 documents, and the largest has 584.
3.3 Non-English-centric translations
For SmolDoc, we additionally collected data for four non-English-centric language pairs, namely from each of the East African languages of Amharic (am) and Swahili (sw) to each of regionally relevant languages Standard Arabic (ar) and Mandarin Chinese (zh). Including the reversed versions of these, this yields 8 total non-English-centric language pairs. Because of the difficulties of choosing/generating good source material in these languages, we used the human translations from English as the source text. However, due to the lack of appropriate evaluation sets, it is difficult to know the value-add of this data over datasets pivoted through English.
4 Data collection and verification
[…]
set | N langs | ex | tok | char | ex/LP | tok/LP | char/LP |
Gatitos | 176 | 693k | 784k | 4.6M | 3.9k | 4.5k | 26k |
SmolSent | 81 | 70k | 994k | 6.1M | 863 | 12k | 75k |
SmolDoc | 100 | 27k | 5.1M | 28M | 263 | 50k | 278k |
[…]
5 Finetuning
To get a general understanding of how helpful these data are, we fine-tune the Gemini 2.0 Flash model (Team et al., 2024a, b) on these data. It is important to note that the experimental set-up in this work is not optimized, and simply evaluates the “out of the box” improvement of simplistic finetuning with no bells and whistles. The ChrF scores could certainly be improved with e.g. a smaller set of languages, mixing in pre-training data, using larger base models, etc..
ChrF on LPs in… | SmolSent (80 LP) | SmolDocSp. (73 LP) | Overlap (38 LP) | |||
Model | 0-shot | 10-shot | 0-shot | 10-shot | 0-shot | 10-shot |
G. Translate | - | - | - | - | 43.2 | - |
NLLB-54b | - | - | - | - | 40.0 | - |
Claude 3.5 Son. | 37.5 | 39.7 | 38.3 | 40.9 | 41.0 | 42.8 |
GPT-4o | 29.9 | 34.1 | 31.8 | 36.3 | 35.4 | 38.5 |
Gemini 2.0 Pro | 38.9 | 38.9 | 39.9 | 40.3 | 42.6 | 42.2 |
[…]
+ SmolSent | 38.3 | 38.3 | 38.8 | 38.8 | 40.6 | 40.6 |
+ SmolDoc | 35.3 | 35.3 | 39.5 | 39.5 | 41.2 | 41.2 |
+ Both | 38.9 | 38.9 | 40.5 | 40.5 | 41.8 | 41.8 |
+ BothG | 39.4 | 39.3 | 41.0 | 40.9 | 42.1 | 42.2 |
[…]
5.1 Evaluation
Since so many language pairs are covered, we evaluate on a combination of all available evaluation sets, namely Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019), and Gatones (Jones et al., 2023). Since no reliable embedding models exist for these languages, we use ChrF with NFKC unicode normalization as our metric.
N-shot: ChrF-counterweighted RAG
To have a strong baseline for N-shot results, we adopt a RAG-based approach that resembles the greedy set-cover algorithm to choose the few-shot exemplars. For each sentence in the eval set, we want the best coverage of the source sentence n-grams as possible, with the least redundancy among exemplars. Therefore, we iteratively choose the exemplar whose source side has the minimum ChrF to the eval source. However, when counting the true positives in the ChrF calculation, we weight the count of each ngram by , where is the number of times has been seen among the exemplars chosen so far, and is a parameter to control how close this algorithm is to ngram set-cover. We use . The set of exemplars we choose from is the concatenation of SmolSent and SmolDocSplit.
5.2 Finetuning Setup and Results
We finetuned Gemini 2.0 Flash for 50 epochs on SmolDoc, SmolSent, a combination of the two (Both), and their combination plus Gatitos (Both+G). To simplify finetuning, we split SmolDoc into sentence pairs (SmolDocSplit). We use a minimalist prompt (Appendix C).
Results can be seen in Table 3. Finetuning on SmolSent gives an average gain of +2.7 ChrF, and SmolDocSplit give the exact same gain on its languages. Mixing the two leads to a gain of +3.3 to +3.6 ChrF, and adding in Gatitos bumps it up to +3.6 to +4.1 ChrF, passing all other 0-shot baselines except Google Translate (on the languages it supports).
The 10-shot RAG results on the un-tuned model are very close to the finetuned 0-shot results, and the finetuned models show no benefit from multi-shot decoding, suggesting that these are two different ways of giving the same information—inference-time versus training time. The 10-shot random results (not included in table) were much lower.
Figure 1 shows the learning curve on a development subset of 37 languages. Although it may be surprising that so many epochs are needed before convergence, we found that further increasing learning rate led to overfitting. The sharp drop near the beginning suggests a domain mismatch between pretraining and finetuning, and suggests that the same data could be used much more effectively with a better training set-up than explored here.
5.3 The Problem with xxen training
Our initial experiments used all data for both enxx and xxen . However, the models lost performance on all tasks. The root cause turned out to be the multiway-parallel data with English targets. Large language models are especially susceptible to repetition in data (Lee et al., 2022), and with 115 language pairs, for every one epoch over the data, the model saw about 115 epochs for each individual target sentence. Therefore, it wildly overfit and lost performance on all language pairs. Mitigating such overfitting is an important research direction to pursue, since many promising datasets are multiway parallel, e.g. Flores-101 (Goyal et al., 2022), Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019), and others. However, this is out of scope for the present paper, so we restrict our experiments to enxx .
Seeing the same source many times likely also has deleterious effects and should also be studied; but these effects, if they exist, are small enough that we were still able to see net gains.
6 Factuality Review
Since SmolDoc contains LLM-generated sources, they contain some factual inaccuracies. We therefore do a full human audit and assign factuality codes to each document. Each of the 584 documents was rated by three raters. Each rating is accompanied by a detailed explanation, including sources cited. Inter-annotator agreement was high, with Cohen’s between each pair of raters between 0.82 and 0.87. The error code distribution can be seen in Fig. 2. Appendix D gives details on the rubric.
All ratings and rationales are made available. In addition, each datum in SmolDoc is given with a simple factuality annotation, which has the value has_errors if any one of the ratings was any of Minor Issues or Clear Issues, and ok otherwise. For some use-cases, like question-answering, practitioners may want to filter out nonfactual data; for others, like translation, one may not be troubled by factual errors. In addition to filtering, this also provides the first factuality dataset for most of these languages.
7 Conclusion
We have open-sourced the Smol dataset, a professionally-translated dataset covering 115 low resource languages and targeting the tasks of translation and factuality. It comprises SmolDoc and SmolSent, two training datasets with the complementary strengths of sentence selection (complex, and high token coverage) and document generation (contextual, varied domains, simpler sentences) respectively. We demonstrate that finetuning Gemini 2.0 Flash on these yields to substantial improvements in translation quality. Smol joins a growing body of resources to support underserved languages in the age of AI.
8 Limitations
The Smol data would benefit from a more thorough review, audit, and correction from community members outside of the translators who created it. Future work on Smol-like datasets should also focus on non-English source text that is not only maximally authentic in the given language, but also covers the topics and concepts most relevant to those languages. This approach is more difficult and would require significant work and review to do correctly. Finally, more research is needed to understand and prevent the overfitting that comes with multi-way parallel data.
9 Ethical implications and Potential Risks
Our released Smol dataset will assist in development and evaluation of machine translation and multilingual AI capabilities for low-resource languages. Although we consider this a net benefit for society, one potential risk is that users may overestimate the quality of machine translations in these languages. Another potential risk is increasing the proliferation of AI-generated content in the target languages on the open web, which may pose issues for future development of web-crawled datasets in these languages. Furthermore, although we have surveyed and talked with native speakers of all Smol languages and found only enthusiasm for technological support of their languages, communities are not monoliths and some community members may prefer that technology not support their language.
Appendix A SmolSent details
A.1 Evaluating the SmolSent selection process
In Section 3.1.1, we describe experiments used to validate the selection process for SmolSent. We train a backbone MT system is pretrained on the Madlad-400 dataset (Kudugunta et al., 2023). The following languages are held out of the training data to be used for fine-tuning experiments: Catalan, Icelandic, Marathi, Turkish, Maltese, Xhosa, Tamil, Basque, and Tajik. The model itself is a 1B parameter encoder-decoder Transformer and is trained from scratch on the Madlad-400 data. Each of the candidate data selection methods is used to select data from the held-out languages, and then each candidate set is used in turn to finetune the backbone model. The results of this finetune step are reported in Table 1, where the set-cover approach is shown to be most effective.
A.2 Notes on Researcher in the Loop
Researcher In the Loop extends the greedy set cover approach thusly: rather than always picking the highest-scoring sentence, we iteratively show the researcher a batch of the 20 highest scoring sentences according to several scores, and let the researcher pick and optionally edit each sentence at each iteration. At each iteration, the researcher may also remove any number of this batch’s sentences from the reservoir. Allowing the researcher to see and edit the sentences allows ensures that the sentences are of high-quality. To deal with the length bias issue, we showed not only sentences that maximize coverage percent, but also that maximize heuristics that weighted the coverage with the number of new tokens hit, like log(coverage_percent)*n_hits .
As described in the paper text, this approach is designed to combat issues such as honeypot sentences. Example Honeypot sentences can be seen in Table 4
Sentence |
Individual determine can get prolonged, reduce along with attractive. |
Sell hand mood situation connect proper decision today spread true. |
Demand indeed off forget act special well treat sometimes notice. |
Agree board book oh trust by attractive supply deal together. |
Picture exactly could ability impact advance then same admire across. |
One physically courage both information language issue laugh common. |
A.3 More detailed corpus statistics on SmolSent
To measure “Bang for our Buck” we define the excess-token ratio as the number distinct tokens in the set cover divided by the number of target tokens, and use it along with the coverage percent to understand the SmolSent dataset. Table 5 compares corpus statistics of SmolSent to four other corpora. sametoks picks a random set of sentences from CommonCrawl until it has the same number of tokens as SmolSent; this only covers 50% of the target tokens and has an excess-token-ratio ratio of 3.3, much worse than SmolSent’s value of 2.3. The samecov baseline randomly picks common-crawl sentences until it has the same token coverage as SmolSent, which necessitates a 67x larger set of sentences, and a correspondingly bloated excess token coverage ratio of 23.1. As a further reference we compare tiers 1 (largest) and 5 (smallest; comparative size to SmolSent) of SmolDoc. As expected of machine-generated text, they have a worse value, corresponding to a narrower spectrum of vocabulary used.
set | N sent | toks | types | cov% | |
SmolSent | 863 | 12k | 5.5k | 2.3 | 99.6% |
sametoks | 863 | 12k | 3.8k | 3.3 | 50.4% |
samecov | 57578 | 877k | 38k | 23.1 | 99.6% |
[…]
SmolDoc-t5 | 820 | 12k | 2.8k | 4.3 | 40.2% |
Appendix B SmolDocDetails
B.1 SmolDoc data tier details
Per-tier statistics on the SmolDoc dataset can be seen in Table 6.
set | N langs | ex | tok | char | ex/LP | tok/LP | char/LP |
SmolDoc-t1 | 5 | 2.9k | 537k | 3.0M | 584 | 107k | 604k |
SmolDoc-t2 | 31 | 14k | 2.7M | 15M | 450 | 88k | 495k |
SmolDoc-t3 | 24 | 6.7k | 1.2M | 6.8M | 280 | 50k | 281k |
SmolDoc-t4 | 8 | 1.0k | 184k | 1.0M | 126 | 23k | 128k |
SmolDoc-t5 | 34 | 2.2k | 401k | 2.2M | 66 | 12k | 65k |
all-Smol | 115 | 97k | 6.1M | 34M | 827 | 52k | 294k |
B.2 Details on SmolDoc prompt creation
To avoid biases from overly tempted prompts, we put in significant effort to make sure the prompts were all very different. Each prompt drew at random from the following elements:
random selection of English words to use in the response
one of 600 manually created topics, e.g. "volcanic eruptions" or "A special tree"
one of 50 tone/tense categories, e.g. "Please use the subjunctive mood.", "Use an effusive tone.""
A style prompt, e.g. "You are the author R.K. Narayan." or "You are a mother talking to her son."
A text modality, e.g. story/dialogue/essay
In addition to this, we added a few more sources of prompts:
Prompts based on urls, meant to simulate different web domains, like Wikipedia and reddit.
Prompts based on continuing the sentences from SmolSent
Prompts based on current events, history, and daily life in different countries
Special effort was made to include dialogues (to get more spoken register) and recipes (unique domain that may also be important to translate).
For each prompt, we generated 8 responses (T=0.7). These were ranked by their simple token density (unique tokens over total tokens), and the top two were chosen for consideration. Using the Researcher-in-the-loop mentality (“measure twice cut once!"), we went over 1000+ responses by hand and scored/edited them. This was mainly to filter out questionable or boring responses. A typical paragraph scored as 0 would be LLM-speak like “X is a complex and multifaceted problem with no easy solution. Here are some suggestions. Keep in mind that there is no one-size-fits all solution, and ultimately, the choice is up to you. […]”.
Example prompts can be seen in Table 7.
Example Prompts |
You are Ernest Hemingway. Write a dialogue about road rage. Use a didactic tone. |
Write a 1-paragraph story concerning an Irish wake. |
You are a teenager talking to his friend. Please carefully craft a 1-paragraph bit about an engineer who subsists off coffee. Try to include the words "confirmed", "move" and "above". |
Give a typical, yet interesting, example of something you would find on reddit. |
Please write a few paragraphs about challenges facing Ethiopia. |
Please write a long passage starting with ’Mum and Dad pause their debate when we hear this creepy clacking that sounds like hail falling.’ |
Write a recipe for baking an almond cake. |
B.3 SmolDoc Errata
Orthographies
Several languages use irregular orthographies. Most notable is Mooré/Mossi (mos), where different translators have used a variety of different conventions. After soliciting community feedback, we plan to release standardized versions to the data.
document selection
When collecting the data for SmolDoc for the Indian languages, we mistakenly included a variety of documents that fell below the corpus diversity threshold described in Section 3.2.
Appendix C Prompts for Decoding
For 0-shot prompting, we used the following, fairly wordy prompt, the SL and TL standing for the source and target language name, respectively:
You are an expert translator. I am going to give you some example pairs of text snippets where the first is in ${SL} and the second is a translation of the first snippet into ${TL}. The sentences will be written ${SL}: <first sentence> ${TL}: <translated first sentence> After the example pairs, I am going to provide another sentence in ${SL} and I want you to translate it into ${TL}. Give only the translation, and no extra commentary, formatting, or chattiness. Translate the text from ${SL} to ${TL}.
For finetuned models, there is no need for such a wordy prompt, and indeed it only risks overfitting. Therefore, we used the following minimalist prompt:
Translate from ${SL} to ${TL}:
Appendix D Factuality rubric
The full factuality rubric given to raters is presented in Table 8.
Rating | Definition of Rating |
N/A - Not Applicable | True/False does not apply here. Most stories, dialogues, or fictional works would be considered N/A, unless they are promoting a falsehood about the real world. |
Not Sure | Claims are made that may not be true, but you aren’t sure. Choose this if it would take over 10 minutes to verify the factuality of the claim. |
No Issues | All claims are factual and accurate. (It’s OK if they are out-of-date, e.g. “Barack Obama is the President of the USA”) |
Minor Issue(s) | There are small inaccuracies. For instance, it may be broadly correct but frame something in a misleading way. |
Clear Issue(s) | There are clear mistakes in factuality. |
Appendix E Translations for NKo
The initial delivery for the NKo language (nqo) had a wide variety of errors. We reached out to the authors from Doumbouya et al. (2023), who did a complete re-translation of the text.
Appendix F Full results
Full per-language results can be seen in Table LABEL:tab:full-results. Results are sorted by the , which is the ChrF of the Both model minus the ChrF of the finetuned Both model—in other words, how much the finetuning on Smol improved the baseline model.
lang | cat | G2F | +sS | +sD | +sB | +sG | Cld | +RAG | G2P | GPT4o | GTr | NLLB | |
ee | Both | +36.1 | 3.0 | 37.7 | 37.5 | 39.1 | 39.2 | 37.8 | 40.0 | 39.5 | 7.5 | 42.7 | 40.7 |
kr | Both | +10.8 | 17.3 | 25.6 | 26.3 | 28.1 | 28.8 | 22.7 | 26.3 | 20.3 | 22.2 | 32.6 | 31.0 |
kg | Both | +9.2 | 34.9 | 46.9 | 35.6 | 44.1 | 43.2 | 43.2 | 47.0 | 37.8 | 29.1 | 50.2 | 3.4 |
bem | Both | +7.3 | 40.0 | 44.8 | 45.1 | 47.3 | 49.2 | 43.3 | 47.7 | 42.3 | 33.3 | 49.7 | 41.8 |
dyu | Both | +5.3 | 17.9 | 22.5 | 22.9 | 23.2 | 23.7 | 23.9 | 24.4 | 21.0 | 4.5 | 22.4 | 12.5 |
din | Both | +4.6 | 20.3 | 23.8 | 23.1 | 24.9 | 25.1 | 23.3 | 25.9 | 21.4 | 1.6 | 25.1 | 26.5 |
st | Both | +4.5 | 49.9 | 54.0 | 55.3 | 54.4 | 55.0 | 49.4 | 57.0 | 53.1 | 49.2 | 49.0 | 47.2 |
luo | Both | +4.1 | 37.4 | 39.1 | 41.3 | 41.5 | 42.0 | 39.1 | 42.0 | 39.6 | 36.1 | 41.3 | 39.5 |
fon | Both | +3.6 | 21.3 | 24.3 | 24.1 | 24.9 | 25.3 | 20.4 | 23.7 | 23.8 | 1.9 | 25.9 | 24.2 |
bm | Both | +3.4 | 30.8 | 28.6 | 34.9 | 34.2 | 34.1 | 34.0 | 36.2 | 33.9 | 9.0 | 35.7 | 32.2 |
ak | Both | +2.6 | 35.5 | 36.1 | 37.6 | 38.1 | 38.2 | 34.4 | 38.1 | 37.3 | 32.2 | 34.5 | 33.3 |
ln | Both | +2.5 | 46.8 | 48.1 | 49.2 | 49.3 | 49.3 | 44.6 | 48.3 | 46.5 | 45.2 | 46.4 | 45.7 |
wo | Both | +1.1 | 30.3 | 30.0 | 30.8 | 31.4 | 31.6 | 31.4 | 32.2 | 30.7 | 29.8 | 36.2 | 30.9 |
tum | Both | +1.1 | 40.8 | 39.5 | 41.9 | 41.9 | 42.8 | 40.0 | 42.7 | 43.8 | 37.7 | 45.4 | 36.2 |
ff | Both | +0.9 | 25.0 | 24.4 | 25.7 | 25.9 | 26.5 | 25.7 | 26.1 | 25.2 | 2.5 | 25.9 | 27.1 |
ti | Both | +0.8 | 24.2 | 24.3 | 24.6 | 25.0 | 25.7 | 25.0 | 26.2 | 26.1 | 9.3 | 26.1 | 25.5 |
yo | Both | +0.6 | 34.6 | 33.8 | 35.4 | 35.2 | 35.8 | 29.2 | 36.8 | 26.7 | 27.6 | 21.3 | 32.6 |
tn | Both | +0.2 | 52.5 | 50.8 | 51.1 | 52.7 | 53.2 | 50.1 | 51.7 | 53.3 | 36.8 | 55.6 | 53.0 |
am | Both | +0.1 | 34.0 | 33.2 | 33.1 | 34.1 | 33.5 | 31.6 | 32.6 | 35.8 | 29.6 | 34.7 | 30.3 |
ig | Both | -0.1 | 47.2 | 47.1 | 47.7 | 47.1 | 47.8 | 43.9 | 46.2 | 47.8 | 46.2 | 47.6 | 46.6 |
so | Both | -0.1 | 49.7 | 46.2 | 50.1 | 49.6 | 49.1 | 48.7 | 49.8 | 50.3 | 50.8 | 50.6 | 48.6 |
sn | Both | -0.5 | 50.5 | 48.3 | 49.9 | 50.0 | 50.3 | 46.8 | 48.8 | 51.8 | 50.3 | 49.2 | 48.2 |
ss | Both | -0.6 | 50.2 | 48.8 | 47.7 | 49.6 | 50.3 | 49.6 | 51.2 | 51.8 | 46.0 | 56.3 | 48.1 |
yue | Both | -0.7 | 26.8 | 25.8 | 25.6 | 26.1 | 26.1 | 28.2 | 28.2 | 27.5 | 31.6 | 25.9 | 22.6 |
om | Both | -0.8 | 40.1 | 38.0 | 39.6 | 39.3 | 39.4 | 39.0 | 40.2 | 41.3 | 38.4 | 41.4 | 39.1 |
lg | Both | -1.1 | 42.5 | 39.9 | 41.0 | 41.4 | 41.7 | 42.0 | 43.1 | 43.5 | 41.0 | 43.6 | 41.1 |
rw | Both | -1.4 | 45.2 | 43.1 | 43.1 | 43.8 | 43.8 | 43.2 | 44.0 | 45.1 | 44.7 | 48.8 | 43.4 |
sw | Both | -1.5 | 66.7 | 64.6 | 64.7 | 65.2 | 64.8 | 64.0 | 65.5 | 67.2 | 66.5 | 65.3 | 60.5 |
mg | Both | -1.9 | 52.8 | 48.9 | 51.1 | 50.9 | 51.5 | 52.4 | 52.5 | 53.3 | 52.2 | 52.6 | 52.1 |
zu | Both | -2.0 | 58.3 | 56.4 | 55.2 | 56.3 | 56.1 | 54.1 | 55.5 | 58.5 | 57.5 | 57.6 | 57.6 |
xh | Both | -2.3 | 53.9 | 49.7 | 51.4 | 51.6 | 51.8 | 51.6 | 52.3 | 53.9 | 53.7 | 54.8 | 51.2 |
nso | Both | -2.3 | 46.8 | 42.8 | 44.0 | 44.5 | 44.6 | 46.9 | 47.7 | 48.1 | 46.9 | 47.6 | 45.5 |
ber | Both | -3.2 | 25.3 | 20.6 | 23.0 | 22.1 | 21.9 | 28.5 | 25.2 | 31.1 | 2.8 | 21.0 | 32.4 |
ha | Both | -3.4 | 54.5 | 50.7 | 50.5 | 51.1 | 51.5 | 50.9 | 51.0 | 54.1 | 53.9 | 53.8 | 53.9 |
ts | Both | -3.6 | 50.6 | 47.2 | 46.5 | 47.0 | 48.1 | 49.7 | 50.1 | 51.6 | 49.0 | 52.9 | 51.3 |
rn | Both | -3.9 | 44.5 | 39.3 | 40.7 | 40.6 | 40.9 | 43.1 | 43.4 | 46.2 | 44.3 | 45.4 | 45.0 |
af | Both | -4.2 | 71.9 | 68.8 | 67.0 | 67.7 | 68.3 | 71.7 | 72.5 | 72.1 | 71.8 | 71.5 | 68.6 |
ny | Both | -5.1 | 55.0 | 47.5 | 50.2 | 49.9 | 49.7 | 53.0 | 53.1 | 55.3 | 53.9 | 55.8 | 50.3 |
[…]
mni-M. | SmolSent | +26.4 | 2.9 | 30.0 | 1.2 | 29.3 | 29.3 | 29.6 | 31.8 | 33.6 | 1.3 | 45.6 | 0.8 |
gaa | Both | +23.1 | 22.7 | 44.5 | 43.7 | 45.8 | 47.4 | 34.7 | 44.0 | 40.9 | 6.6 | 48.3 | - |
dov | Both | +21.1 | 19.1 | 39.2 | 39.4 | 40.2 | 40.6 | 19.2 | 39.5 | 18.2 | 8.7 | 41.7 | - |
ahr | SmolDoc | +17.8 | 24.2 | 31.8 | 41.5 | 42.0 | 42.8 | 32.8 | 39.0 | 30.0 | 36.9 | - | - |
sus | Both | +17.8 | 11.3 | 28.3 | 27.0 | 29.1 | 30.3 | 26.1 | 29.4 | 20.7 | 5.6 | 34.6 | - |
nqo | Both | +17.5 | 0.2 | 17.9 | 17.3 | 17.7 | 17.5 | 17.1 | 17.9 | 17.2 | 1.1 | 19.1 | - |
alz | Both | +15.5 | 16.9 | 31.5 | 31.7 | 32.4 | 33.4 | 25.3 | 30.6 | 26.9 | 8.9 | 36.6 | - |
lu | Both | +13.8 | 27.6 | 37.5 | 39.0 | 41.4 | 42.2 | 27.2 | 37.0 | 34.8 | 21.9 | - | - |
cgg | Both | +12.2 | 32.6 | 40.5 | 40.5 | 44.8 | 44.8 | 37.3 | 42.2 | 37.7 | 28.3 | 42.8 | - |
ks-D. | SmolSent | +11.7 | 14.9 | 26.6 | 18.7 | 26.6 | 27.7 | 23.8 | 27.1 | 21.5 | 19.2 | - | 21.1 |
brx | SmolSent | +11.6 | 24.3 | 36.0 | 0.4 | 35.9 | 37.0 | 30.8 | 35.9 | 36.2 | 5.2 | - | - |
mag | SmolDoc | +8.1 | 47.0 | 45.3 | 55.4 | 55.1 | 54.7 | 47.3 | 51.8 | 47.4 | 48.7 | - | 59.4 |
ki | Both | +7.7 | 32.6 | 38.2 | 38.8 | 40.3 | 40.6 | 35.9 | 42.0 | 39.1 | 10.5 | - | 38.4 |
aa | Both | +7.4 | 14.2 | 20.1 | 20.3 | 21.6 | 21.8 | 18.9 | 20.6 | 18.9 | 5.6 | 23.1 | - |
ks | SmolSent | +7.3 | 22.1 | 29.4 | 0.4 | 29.4 | 29.7 | 28.0 | 30.4 | 30.5 | 26.3 | - | 36.7 |
nr | Both | +7.0 | 48.9 | 54.0 | 50.2 | 55.9 | 57.5 | 48.0 | 53.6 | 51.2 | 45.5 | 59.5 | - |
doi | SmolDoc | +6.6 | 34.3 | 28.1 | 40.8 | 40.9 | 41.3 | 35.9 | 39.5 | 38.2 | 27.7 | 40.4 | - |
sat-L. | SmolSent | +6.4 | 12.8 | 19.5 | 15.9 | 19.2 | 20.8 | 25.3 | 22.5 | 22.7 | 21.3 | 22.7 | - |
mfe | Both | +5.3 | 59.5 | 65.4 | 61.7 | 64.8 | 66.9 | 59.6 | 65.0 | 59.8 | 59.5 | 67.5 | - |
ach | Both | +5.2 | 33.2 | 43.1 | 32.3 | 38.4 | 39.2 | 32.4 | 37.3 | 35.1 | 23.8 | 43.2 | - |
ayl | Both | +4.5 | 47.3 | 51.7 | 51.3 | 51.8 | 53.9 | 45.3 | 48.9 | 46.3 | 48.6 | - | - |
ber-L. | Both | +4.2 | 26.1 | 27.9 | 29.7 | 30.3 | 30.9 | 27.6 | 32.7 | 32.1 | 21.2 | 34.7 | - |
apd-S. | Both | +3.6 | 42.3 | 50.2 | 43.8 | 45.9 | 47.1 | 43.2 | 45.0 | 42.5 | 45.6 | - | - |
ve | Both | +3.5 | 47.9 | 50.0 | 49.5 | 51.4 | 52.2 | 50.2 | 53.1 | 52.7 | 43.9 | 56.8 | - |
kri | Both | +2.8 | 31.5 | 34.2 | 32.1 | 34.3 | 34.7 | 34.5 | 33.5 | 30.7 | 34.9 | 34.9 | - |
tiv | Both | +2.5 | 23.8 | 25.7 | 25.5 | 26.3 | 26.5 | 22.3 | 24.2 | 24.5 | 1.5 | 25.2 | - |
gn | SmolSent | +2.3 | 37.4 | 37.8 | 29.4 | 39.7 | 38.4 | 36.0 | 38.0 | 36.4 | 35.6 | 38.4 | 38.5 |
mos | Both | +2.3 | 18.2 | 20.9 | 19.2 | 20.5 | 21.1 | 24.3 | 25.0 | 20.9 | 1.3 | - | 23.8 |
ar-M. | Both | +0.1 | 40.1 | 41.8 | 38.7 | 40.2 | 40.9 | 40.5 | 40.8 | 40.4 | 41.0 | - | 43.0 |
arz | Both | -0.1 | 48.6 | 46.1 | 48.6 | 48.5 | 47.8 | 49.6 | 50.3 | 48.8 | 49.7 | - | 49.6 |
kl | SmolSent | -0.3 | 40.6 | 39.7 | 29.0 | 40.3 | 41.2 | 42.2 | 43.1 | 41.2 | 41.6 | 42.9 | - |
sa | SmolSent | -0.3 | 33.0 | 32.9 | 27.4 | 32.7 | 33.4 | 31.8 | 33.0 | 32.1 | 32.0 | 35.2 | 29.0 |
ay | SmolSent | -0.5 | 32.7 | 31.8 | 23.2 | 32.2 | 32.4 | 33.4 | 33.2 | 32.9 | 30.0 | 34.7 | 31.7 |
efi | Both | -0.6 | 14.7 | 14.5 | 14.3 | 14.1 | 14.2 | 15.3 | 15.1 | 15.0 | 2.2 | - | - |
bci | Both | -0.7 | 23.2 | 22.2 | 22.4 | 22.5 | 22.9 | 17.1 | 20.7 | 27.6 | 1.0 | 29.8 | - |
ndc-Z. | Both | -1.0 | 29.2 | 27.9 | 27.8 | 28.2 | 28.3 | 27.6 | 28.0 | 29.6 | 28.6 | 29.5 | - |
es | SmolSent | -1.1 | 62.4 | 61.3 | 51.3 | 61.3 | 61.3 | - | - | 63.0 | - | 63.5 | 61.8 |
sat | SmolSent | -1.3 | 32.4 | 30.9 | 0.6 | 31.1 | 30.8 | 34.7 | 36.0 | 36.3 | 1.8 | 35.7 | - |
nd | Both | -1.4 | 43.9 | 41.5 | 42.5 | 42.5 | 43.2 | 42.3 | 42.9 | 44.5 | 43.6 | - | - |
qu | SmolSent | -1.9 | 34.7 | 32.8 | 29.3 | 32.8 | 33.0 | 35.3 | 35.1 | 34.0 | 22.0 | 36.3 | 27.9 |
lus | SmolSent | -2.1 | 42.6 | 39.7 | 37.2 | 40.5 | 41.4 | 40.6 | 41.5 | 43.8 | 33.5 | 42.6 | 39.0 |
scn | SmolDoc | -2.2 | 52.4 | 47.3 | 51.2 | 50.2 | 50.8 | 49.9 | 51.4 | 52.1 | 49.5 | 53.3 | 51.0 |
ne | SmolDoc | -2.7 | 54.3 | 51.8 | 51.9 | 51.6 | 52.1 | 52.4 | 52.5 | 52.4 | 52.7 | 54.9 | 45.2 |
pa-A. | SmolSent | -3.0 | 38.1 | 35.8 | 0.3 | 35.1 | 35.7 | 41.6 | 41.2 | 36.7 | 37.3 | 43.5 | - |
aeb | Both | -3.3 | 46.5 | 41.9 | 42.7 | 43.2 | 43.4 | 45.9 | 46.8 | 47.6 | 49.2 | - | 43.8 |
bo | SmolSent | -4.3 | 41.3 | 36.7 | 34.5 | 37.0 | 37.3 | 42.6 | 42.1 | 43.3 | 19.8 | 41.8 | 36.9 |
pcm | Both | -6.5 | 47.9 | 43.5 | 39.7 | 41.4 | 41.6 | 51.3 | 45.7 | 49.8 | 56.0 | - | - |
tcy | SmolDoc | -6.8 | 34.7 | 22.0 | 28.0 | 27.9 | 28.8 | 28.2 | 29.3 | 36.7 | 21.6 | 39.1 | - |
ktu | Both | -9.4 | 56.6 | 59.3 | 40.1 | 47.2 | 51.3 | 45.8 | 48.4 | 57.8 | 22.3 | 64.3 | - |
Appendix G Complete Per-Language details: the Big-Smol table
A summary of all Smol language pairs and coarse-grained information about them can be seen in Table LABEL:tab:full-details. Numbers are given in terms of examples; keep in mind that a single example in SmolDoc is a document, whereas in SmolSent it is a sentence.
Lang. pair | target language name | ISO 15924 Script | Continent | SmolSent ex | SmolDoc ex |
am_ar | Arabic | Arab | Asia | 0 | 329 |
am_zh | Mandarin Chinese | Hans | Asia | 0 | 329 |
en_aa | Afar | Latn | Africa | 863 | 66 |
en_ach | Acoli | Latn | Africa | 863 | 66 |
en_aeb | Tunisian Arabic | Arab | Africa | 862 | 66 |
en_af | Afrikaans | Latn | Africa | 863 | 130 |
en_ahr | Ahirani | Deva | Asia | 0 | 457 |
en_ak | Twi | Latn | Africa | 863 | 260 |
en_alz | Alur | Latn | Africa | 863 | 66 |
en_am | Amharic | Ethi | Africa | 863 | 584 |
en_apd | Sudanese Arabic | Arab | Africa | 855 | 66 |
en_ar-MA | Morrocan Arabic | Arab | Africa | 863 | 260 |
en_arz | Egyptian Arabic | Arab | Africa | 863 | 260 |
en_ay | Aymara | Latn | Americas | 863 | 0 |
en_ayl | Libyan Arabic | Arab | Africa | 863 | 66 |
en_bci | Baoulé | Latn | Africa | 863 | 66 |
en_bem | Bemba (Zambia) | Latn | Africa | 863 | 66 |
en_ber | Tamazight (Tifinagh Script) | Tfng | Africa | 862 | 130 |
en_ber-Latn | Tamazight (Latin Script) | Latn | Africa | 862 | 130 |
en_bfq | Badaga | Taml | Asia | 0 | 457 |
en_bfy | Bagheli | Deva | Asia | 0 | 457 |
en_bgq | Bagri | Deva | Asia | 0 | 457 |
en_bm | Bambara | Latn | Africa | 863 | 260 |
en_bns | Bundeli | Deva | Asia | 0 | 456 |
en_bo | Tibetan | Tibt | Asia | 863 | 0 |
en_bra | Braj | Deva | Asia | 0 | 457 |
en_brx | Bodo (India) | Deva | Asia | 863 | 0 |
en_ccp-Latn | Chakma (Latin script) | Latn | Asia | 0 | 457 |
en_cgg | Chiga | Latn | Africa | 863 | 66 |
en_dhd | Dhundari | Deva | Asia | 0 | 456 |
en_din | Dinka | Latn | Africa | 863 | 66 |
en_doi | Dogri | Deva | Asia | 0 | 454 |
en_dov | Dombe | Latn | Africa | 863 | 66 |
en_dyu | Dyula | Latn | Africa | 863 | 66 |
en_ee | Ewe | Latn | Africa | 863 | 130 |
en_efi | Efik | Latn | Africa | 863 | 66 |
en_ff | Fulfulde | Latn | Africa | 862 | 260 |
en_fon | Fon | Latn | Africa | 863 | 66 |
en_gaa | Ga | Latn | Africa | 863 | 66 |
en_gn | Guarani | Latn | Americas | 863 | 0 |
en_grt-Latn | Garo (Latin script) | Latn | Asia | 0 | 457 |
en_ha | Hausa | Latn | Africa | 863 | 584 |
en_hoc-Wara | Ho (Warang Chiti script) | Wara | Asia | 0 | 457 |
en_ig | Igbo | Latn | Africa | 863 | 391 |
en_kfy | Kumaoni | Deva | Asia | 0 | 457 |
en_kg | Kongo | Latn | Africa | 863 | 66 |
en_ki | Kikuyu | Latn | Africa | 863 | 66 |
en_kl | Kalaallisut | Latn | Americas | 863 | 0 |
en_kr | Kanuri | Latn | Africa | 863 | 66 |
en_kri | Krio | Latn | Africa | 863 | 130 |
en_kru | Kurukh | Deva | Asia | 0 | 457 |
en_ks | Kashmiri | Arab | Asia | 863 | 0 |
en_ks-Deva | Kashmiri (Devanagari script) | Deva | Asia | 863 | 0 |
en_ktu | Kituba (DRC) | Latn | Africa | 863 | 66 |
en_lep | Lepcha | Lepc | Asia | 0 | 456 |
en_lg | Luganda | Latn | Africa | 863 | 260 |
en_lif-Limb | Limbu (Limbu script) | Limb | Asia | 0 | 457 |
en_ln | Lingala | Latn | Africa | 863 | 260 |
en_lu | Kiluba (Luba-Katanga) | Latn | Africa | 863 | 66 |
en_luo | Luo | Latn | Africa | 863 | 260 |
en_lus | Mizo | Latn | Asia | 863 | 0 |
en_mag | Magahi | Deva | Asia | 0 | 456 |
en_mfe | Morisien | Latn | Africa | 863 | 66 |
en_mg | Malagasy | Latn | Africa | 863 | 391 |
en_mjl | Mandeali | Deva | Asia | 0 | 457 |
en_mni-Mtei | Meiteilon (Manipuri) | Mtei | Asia | 863 | 0 |
en_mos | Mossi | Latn | Africa | 863 | 66 |
en_mtr | Mewari | Deva | Asia | 0 | 457 |
en_nd | North Ndebele | Latn | Africa | 863 | 66 |
en_ndc-ZW | Ndau | Latn | Africa | 863 | 66 |
en_ne | Nepali | Deva | Asia | 0 | 456 |
en_noe | Nimadi | Deva | Asia | 0 | 315 |
en_nr | South Ndebele | Latn | Africa | 863 | 66 |
en_nso | Sepedi | Latn | Africa | 863 | 130 |
en_ny | Chichewa | Latn | Africa | 863 | 260 |
en_nqo | NKo | Nkoo | Africa | 863 | 66 |
en_om | Oromo | Latn | Africa | 863 | 391 |
en_pa-Arab | Lahnda Punjabi (Pakistan) | Arab | Asia | 863 | 0 |
en_pcm | Nigerian Pidgin | Latn | Africa | 863 | 130 |
en_qu | Quechua | Latn | Americas | 863 | 0 |
en_rn | Rundi | Latn | Africa | 863 | 66 |
en_rw | Kinyarwanda | Latn | Africa | 863 | 260 |
en_sa | Sanskrit | Deva | Asia | 863 | 0 |
en_sat | Santali (Ol Chiki script) | Olck | Asia | 863 | 0 |
en_sat-Latn | Santali (Latin Script) | Latn | Asia | 863 | 0 |
en_scl | Shina | Arab | Asia | 0 | 457 |
en_scn | Sicilian | Latn | Europe | 0 | 100 |
en_sd-Deva | Sindhi (Devanagari script) | Deva | Asia | 0 | 456 |
en_sgj | Surgujia | Deva | Asia | 0 | 356 |
en_sjp | Surjapuri | Deva | Asia | 0 | 299 |
en_sn | Shona | Latn | Africa | 863 | 260 |
en_so | Somali | Latn | Africa | 862 | 260 |
en_spv | Sambalpuri | Orya | Asia | 0 | 457 |
en_ss | Swati | Latn | Africa | 863 | 66 |
en_st | Sesotho | Latn | Africa | 863 | 260 |
en_sus | Susu | Latn | Africa | 863 | 66 |
en_sw | Swahili | Latn | Africa | 863 | 584 |
en_tcy | Tulu | Knda | Asia | 0 | 451 |
en_ti | Tigrinya | Ethi | Africa | 863 | 260 |
en_tiv | Tiv | Latn | Africa | 863 | 66 |
en_tn | Tswana | Latn | Africa | 863 | 66 |
en_trp | Kok Borok | Latn | Asia | 0 | 457 |
en_ts | Tsonga | Latn | Africa | 863 | 66 |
en_tum | Tumbuka | Latn | Africa | 863 | 66 |
en_unr-Deva | Mundari (Devanagari script) | Deva | Asia | 0 | 457 |
en_ve | Venda | Latn | Africa | 863 | 66 |
en_wbr | Wagdi | Deva | Asia | 0 | 455 |
en_wo | Wolof | Latn | Africa | 863 | 260 |
en_xh | Xhosa | Latn | Africa | 863 | 260 |
en_xnr | Kangri | Deva | Asia | 0 | 457 |
en_xsr-Tibt | Sherpa (Tibetan script) | Tibt | Asia | 0 | 457 |
en_yo | Yoruba | Latn | Africa | 863 | 584 |
en_yue | Cantonese | Hant | Asia | 863 | 584 |
en_zu | Zulu | Latn | Africa | 863 | 260 |
sw_ar | Arabic | Arab | Asia | 0 | 330 |
sw_zh | Mandarin Chinese | Hans | Asia | 0 | 330 |