### v2

# Professionally translated parallel data for 115 under-represented languages

Abstract
We open-source Smol (Set of Maximal Overall Leverage),111
google/smol a suite of
training data
to unlock machine translation for low-resource languages. Smol has been translated into 124 (and growing) under-resourced languages (125 language pairs),222Experiments are mostly on a subset of 115 languages, before volunteer translations of additional languages finished. The paper title reflects this. including many for which there exist no previous public resources, for a total of 6.1M translated tokens. Smol comprises two sub-datasets, each carefully chosen for maximum impact given its size: SmolSent, a set of sentences chosen for broad unique token coverage, and SmolDoc, a document-level resource focusing on a broad topic coverage.
They join the already released Gatitos for a trifecta of paragraph, sentence, and token-level content.
We demonstrate that using Smol to prompt or fine-tune Large Language Models yields robust ChrF improvements. In addition to translation, we provide factuality ratings and rationales for all documents in SmolDoc, yielding the first factuality datasets for most of these languages.
SMOL:
Professionally translated parallel data
for 115 under-represented languages
Isaac Caswell∗ Elizabeth Nielsen∗ Jiaming Luo Colin Cherry Geza Kovacs Hadar Shemtov Partha Talukdar Dinesh Tewari Baba Mamadi Diane Djibrila Diane Solo Farabado Cissé Koulako Moussa Doumbouya Edoardo Ferrante Alessandro Guasoni Christopher Homan Mamadou K. Keita Sudhamoy DebBarma Ali Kuzhuget David Anugraha Muhammad Ravi Shulthan Habibi Genta Indra Winata Anthony Munthali Sina Ahmadi Andrei Chemyshev Mingfei Lau Jonathan Eng Google {Research, Deepmind} NKo USA INC Stanford University Conseggio pe-o patrimònio linguistico ligure Universitas Indonesia University of Zurich Rochester Institute of Technology Capital One tyvan.ru Mari Institute of Language, Literature and History Led volunteer contributions for NKo, Ligurian, Zarma, Mooré, Kokborok, Russian, Indonesian, Hawrami+Gilaki+Sorani, Tumbuka, Meadow Mari, and Cantonese, respectively ∗{icaswell, eknielsen}@google.com
1 Introduction
There exist no professionally-translated data for most of the world’s 7000 or so languages, rendering tasks like Machine Translation near impossible. High-quality data is needed. However, it is not clear how best to use a limited budget for an expensive task like professional translation. As shown by the Gatitos dataset (Jones et al., 2023), word-level translations provide large benefits to translation quality for low-resource languages at the lowest cost. However, gains quickly saturate, as single tokens are not very expressive. Sentence-level data is better for a model once token-level data saturates, but it has much more inherent redundancy; and document-level data is even more effective…and more redundant.
In this work, we release the Smol dataset, which provides professionally translated sentence- and document-level data for 124 LRLs (125 language pairs). Smol contains two sub-datasets:
SmolSent: 863 English sentences covering 5.5k of the most common English tokens,333In this paper, ‘token’ refers to typographic units as an approximation to words, not subword tokens from a model’s vocabulary. professionally translated into 88 languages.
SmolDoc 584 English documents covering a wide range of topics, domains, and tokens, generated by an LLM and professionally translated into 106 languages.
We demonstrate the utility of these data for finetuning and prompting LLMs for translation, and provide factuality annotations for all documents.
2 Related work
There are not many training datasets with human-translated data for Low-Resource Languages (LRLs), where we operationally define LRL as any language beyond the first 100 supported by most traditional crawls and MT providers (enumerated in Appendix section A).
Tatoeba (Tiedemann, 2020) is probably the most multilingual, but it is made of volunteer contributions and of unclear quality. The Gatitos dataset Jones et al. (2023) consists of a 4000-entry lexicon translated into 170 LRLs, but is only token-level. Most similar to the present work, NLLB-Seed is a high-quality, sentence-level training set of 6k sentences selected from English Wikipedia and professionally translated into 44 LRLs Team et al. (2022). There are also several professionally-translated evaluation sets, namely Flores-101 and Flores-200 Goyal et al. (2022); Team et al. (2022), and NTreX (Federmann et al., 2022a).
While highly multilingual, professionally translated training data is rare, there is a growing number of bottom-up community data sources organized through research collectives like Masakhane ( et al., 2020), Turkish Interlingua (Mirzakhalov et al., 2021a, b), and GhanaNLP (Azunre et al., 2021a); and conferences and workshops like AfricaNLP, AmericasNLP (Mager et al., 2021) and ArabNLP. These datasets are usually generated by researchers fluent in the languages, and are therefore especially high quality. In addition to providing datasets, such efforts frequently also provide models and baselines, or even public interfaces, like the Khaya Translator Web App444https://ghananlp.org/project/translator-webapp/ by GhanaNLP for West African languages, and the lesan.ai555https://lesan.ai/translate translation website for Ethiopian languages.
Participation is especially strong from the African continent, including corpora and models for pan-East-African languages (Babirye et al., 2022), languages from the Horn of Africa (Hadgu et al., 2022), Ethiopian languages (Teferra Abate et al., 2018; Gezmu et al., 2021), Ugandan languages (Akera et al., 2022), Emakhuwa (Ali et al., 2021), South-African languages (Eiselen and Puttkammer, 2014), Setswana and Sepedi (Marivate et al., 2020), Yorùbá (Adelani et al., 2021b, a), Oshiwambo (Nekoto et al., 2022), Igbo (Ezeani et al., 2020), Zulu (Mabuya et al., 2021), Twi (Azunre et al., 2021b), Gbe (Hacheme, 2021), Bambara (Tapo et al., 2021), and Fon (Emezue and Dossou, 2020). Outside of Africa, corpora have been created for languages of the Americas, including for four indigenous languages of Peru in Bustamante et al. (2020), the numerous results on the largely South- and Central American languages from the first AmericasNLP conference (Mager et al., 2021), and the Inuktitut language of Canada (Joanis et al., 2020). Datasets for lower-resourced languages of India have also sprung up, including the 13-language PMIndia (Haddow and Kirefu, 2020), and datasets focused on languages of the Northeast like Mizo (Thihlum et al., 2020), Khasi (Laskar et al., 2021) and Assamese (Laskar et al., 2020). Further West, PARME (Ahmadi et al., 2025) has provided some of the first human-translated content for Kurdish and Iranian languages. Finally, a variety of such datasets and models are available for public use on HuggingFace666https://huggingface.co/datasets?multilinguality=multilinguality:translation&task_categories=task_categories:translation or Zenodo.777https://zenodo.org/communities/africanlp/
3 Text Selection
Translation requires significant investment and can’t be easily re-done, so great care needs be put into carefully choosing sentences to translate. For both sub-datasets SmolDoc and SmolSent, selection or generation of source text is done in English. Selecting only English has clear biases, but also has advantages—most notably, for N languages, it requires N times less work to quality control. Future work should consider focusing on non-English sources.
3.1 SmolSent: Token Set Cover
Our basic motivation for creating SmolSent was to help models overcome vocabulary issues, which are common for the lowest-resource languages (Nielsen et al., 2025; Bapna et al., 2022). Therefore, we frame this as a set-cover problem, and pick the smallest set of sentences (from CommonCrawl888https://commoncrawl.org/ we use all available snapshots as of August 20, 2022) that covers the largest set of target tokens. The tokens we chose to cover (the target set) were the English side of Gatitos, as well as the most common 2,500 tokens from an English web crawl. Set cover is NP-hard, so we approximate it with a greedy algorithm that iteratively picks the sentence with the highest coverage percent, defined as the percentage of its tokens that are in the target set.
Preliminary work on Token Set-Cover
| Method | ChrF |
| Random | 30.5 |
| Token set-cover | 31.7 |
| N-gram DWD | 30.0 |
| Embedding DWD | 27.5 |
To evaluate the token set-cover approach, we started by selecting data from existing web-scraped parallel data. We pretrain a multilingual Neural Machine Translation (NMT) model on parallel data from 294 language pairs from Madlad-400 Kudugunta et al. (2023), with nine languages held out to simulate LRLs. We fine-tune this model on sets of existing parallel data in each of the held-out languages, and evaluate on Flores-200. Details on the experimental set-up in Appendix B.1.
In addition to Greedy Token Set-Cover, we explore two methods that balance data diversity and data quality. First, we implement Ambati et al. (2011)’s ‘density-weighted diversity’ (DWD) metric, which is an -gram based metric for diversity and quality. Second, we implement an embedding-based version of DWD, which takes the weighted harmonic mean of perplexity under the Palm 2 model Anil et al. (2023) (proxy for quality), and embedding distance on mBERT sentence embeddings (proxy for diversity). We apply both methods to the English side of the parallel data only, to simulate the case where we don’t yet have LRL translations. As a baseline, we randomly select sentences.
Table 1 shows results after finetuning. Greedy token set-cover performs the best, with diversity-based metrics actively hurting performance.
Researcher in the Loop (RITL)
Despite its success in the ablation, Greedy Token Set Cover had several problems when we scaled it to select from among all the English sentences of CommonCrawl. Firstly, it is maximized by honeypots, or nonsense strings dense in content words (Appendix Table B.1); and secondly, it biases towards short sentences, causing length distribution artifacts.
These problems are not easy to solve with heuristics—for example, if you disqualify lists with commas you’ll get ones with spaces, if you require sentences to have some function words or token-length diversity, you’ll get other sorts of garbled sentences, and so on. However, a dataset like Smol is small enough to manually inspect. Therefore we develop Researcher in the Loop Greedy Set-Cover (Algorithm 1), where the domain expert (the researcher) can look at and edit each individual sentence.999This work was conducted before the advent of LLMs. Today, this could be simplified using LLMs as autoraters. The result of this process is SmolSent, a set which uses 863 sentences to cover 5519 unique tokens. Qualitatively, SmolSent consists of complex sentences with wide vocabulary coverage; quantitative metrics are explored in Appendix B.3.
3.2 SmolDoc: LLMs with prompt mesh
SmolDoc follows a different and complementary approach. Whereas SmolSent consists of a small set of sentences that are selected from natural text, are complex, and cover many tokens; SmolDoc instead consists of documents that are generated and are simpler, but cover many topics. It should be noted that the token-coverage approach described above failed resoundingly for longer documents, as the prevalence of the honeypots was magnified.
To generate SmolDoc, we used a collection of templates to create a few thousand diverse prompts with a wide range of topics, domains, words, tenses, grammatical cases, and registers (e.g. formal, informal). Appendix C.2 gives details and examples.
Corpus Diversity Ranking for SmolDoc
Document-by-document evaluation as described above does not help one understand corpus diversity—for example, if an almost identical document appears twice, only one of them should be included. Therefore, we rank all candidates by how much new information they add to the corpus, by iteratively finding the document contributing the least new information and removing it, thus ranking all documents. Our criterion for “new information” was the average character 9-gram Inverse Document Frequency (IDF) score of a document—in other words, how rare its substrings were across all of the documents in the pool so far. To down-weight internally repetitive documents, we substracted the fourth moment Bread score (Caswell et al., 2023).
Language Tiers for SmolDoc
We wanted to translate more data for languages with more speakers. We break the languages into the five different groups, each with a larger subset of the generated documents. Each tier contains translations of the top N documents as ranked by corpus diversity. These can be seen in Appendix Table C.1.
Non-English-centric translations
For SmolDoc, we additionally collected data for four non-English-centric language pairs, from each of the East African languages of Amharic (am) and Swahili (sw) to each of regionally relevant languages Standard Arabic (ar) and Mandarin Chinese (zh). Including the reversed versions of these, this yields 8 total language pairs. Because of the difficulty of generating good source material in these languages, we used the existing SmolDoc translations to Swahili/Amharic as the source text. However, due to the lack of appropriate evaluation sets, it is hard to know the value-add of this data over datasets pivoted through English.
4 Data Collection and Verification
Several languages are contributed by volunteers; they are listed as co-authors.101010Community contributions of translations or corrections are welcome; please reach out to the authors or join the TUSL Discord. For the other languages, the translation provider we contracted has worked with us for many years, and has a pre-existing relationship with professional translators for all languages in the Smol datasets. The translators are paid a fair wage, and their identities are contractually kept anonymous to us. We checked the delivery for duplicate translations, anomalous source/target length ratios, and similarity with Google Translate outputs. Very few languages were flagged this way. Following this, we ran FunLangID (Caswell, 2024) on all segments and found no issues. Manual inspection turned up several issues with nonunicode fonts (e.g. ô for =) for West African languages, and nonstandard orthography for Santali; these issues were then fixed. The choice of script, orthography and translation variety was challenging for many communities, including Kurdish, Zaza-Gorani and Gilaki languages, all of which have more than one orthography and lack a standard variety.
The largest missing check is for fluency, which is hard to measure without trusted native speakers outside of the translation agency, or trusted LLMs; neither of which exist for all Smol languages.
| Total Dataset | Per Language Pair (LP) | ||||||
| Set | # languages | Examples | Tokens | Characters | Examples | Tokens | Characters |
| Gatitos | 178 | 714k | 807k | 4.8M | 4.0k | 4.5k | 27k |
| SmolSent | 88 | 76k | 1.1M | 6.6M | 863 | 12k | 75k |
| SmolDoc | 106 | 27k | 5.1M | 29M | 256 | 48k | 270k |
| Smol{Sent+Doc} | 125 | 103k | 6.2M | 35M | 824 | 50k | 282k |
| LP subset | Smol-Sent (80 LP) | Smol-Doc (73 LP) | Intersect (38 LP) | Hard (32 LP) | ||||
| Model | 0-shot | 10-shot | 0-shot | 10-shot | 0-shot | 10-shot | 0-shot | 10-shot |
| G. Translate | - | - | - | - | 43.2 | - | - | - |
| NLLB-54b | - | - | - | - | 40.0 | - | - | - |
| Claude 3.5 Son. | 37.5 | 39.7 | 38.3 | 40.9 | 41.0 | 42.8 | 30.0 | 33.5 |
| GPT-4o | 29.9 | 34.1 | 31.8 | 36.3 | 35.4 | 38.5 | 15.9 | 23.7 |
| Gemini 2.0 Pro | 38.9 | 38.9 | 39.9 | 40.3 | 42.6 | 42.2 | 31.4 | 31.7 |
| Gemini-2.0 Flash | 35.6 | 38.4 | 36.9 | 39.7 | 40.2 | 41.4 | 26.3 | 30.4 |
| + SmolSent | 38.3 | 38.3 | 38.8 | 38.8 | 40.6 | 40.6 | 32.5 | 32.6 |
| + SmolDoc | 35.3 | 35.4 | 39.5 | 39.5 | 41.2 | 41.2 | 31.8 | 31.8 |
| + Both | 38.9 | 38.9 | 40.5 | 40.5 | 41.8 | 41.8 | 33.4 | 33.4 |
| + Both+G | 39.4 | 39.3 | 41.0 | 40.9 | 42.1 | 42.2 | 33.9 | 33.9 |
| +3.8 | +0.9 | +4.1 | +1.2 | +1.9 | +0.8 | +7.6 | +3.5 | |
5 Finetuning and In-Context Learning
We use fine-tuning and ICL as tools to demonstrate the value of the Smol dataset. As this is a data paper, these experiments are motivated by the maxim “what could any researcher simply train with public APIs?” More involved techniques, e.g. Reinforcement-Learning-based approaches, will likely lead to stronger results.
5.1 Evaluation
Since so many language pairs are covered, we evaluate on a combination of all available evaluation sets, namely Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019),
and an in-house eval set. Since no reliable embedding models exist for these languages, trained metrics are not an option, so we use ChrF Popović (2015) as implemented in SacreBleu (Post, 2018)111111signature: case.mixed+numchars.6+numrefs.1
+space.False+tok.none+version.1.3.0 with NFKC unicode normalization as our metric. For ten-shot decoding, exemplars were selected from both sub-datasets of Smol using ChrF-counterweighted RAG (Appendix D).
5.2 Finetuning Setup and Results
We finetuned Gemini 2.0 Flash for 40 epochs on SmolDoc, SmolSent, a combination of the two (Both), and their combination plus Gatitos (Both+G). To simplify finetuning, we split SmolDoc into sentence pairs (SmolDocSplit).
Results can be seen in Table 3. Finetuning on SmolSent gives an average gain of +2.7 ChrF points, and SmolDocSplit gives +2.6 ChrF points on its languages. Concatenating the two training datasets leads to a gain of +3.3 to +3.6 ChrF points, and adding in Gatitos bumps it up to +3.8 to +4.1 ChrF points, passing all baselines except Google Translate. The 10-shot RAG results on the un-tuned model are very close to the finetuned 0-shot results, and the finetuned models show no benefit from multi-shot decoding, suggesting that these are two different ways of giving the same information—inference-time versus training time. The 10-shot random results (not included in table) were much lower.
Gains were highest on languages that are not related to mid- or high-resource languages, and lowest on dialects close to major languages. As a heuristic to measure this, we exclude languages that are on Google Translate or closely related to languages on it (Appendix G). The average gain on these languages jumps to +7.6 ChrF.
Figure 1 shows the learning curve on a development subset of 37 languages. Although it may be surprising that so many epochs are needed before convergence, we found that further increasing learning rate led to overfitting. The sharp drop near the beginning suggests a domain mismatch between pretraining and finetuning, and suggests that the same data could be used much more effectively with a better training set-up than explored here.
5.3 The Problem with xxen training
Our initial experiments used all data for both enxx and xxen . However, the models lost performance on all tasks. The root cause turned out to be the multiway-parallel data with English targets. LLMs are especially susceptible to repetition in data (Lee et al., 2022), and with 115 language pairs, for every one epoch over the data, the model saw about 115 epochs for each individual target sentence. Therefore, it wildly overfit and lost performance on all language pairs. Mitigating such overfitting is an important research direction to pursue, since many promising datasets are multiway parallel, e.g. Flores-101 (Goyal et al., 2022), Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019), and others. However, this is out of scope for the present paper, so we restrict our experiments to enxx .
Seeing the same source many times likely also has deleterious effects and should also be studied; but these effects, if they exist, are small enough that we were still able to see net gains.
6 Factuality Review
Since SmolDoc contains LLM-generated sources, they contain some factual inaccuracies. We therefore do a full human audit and assign factuality codes to each document. Each of the 584 documents was rated by three raters. Each rating is accompanied by a detailed explanation, including sources cited. Inter-annotator agreement was high, with Cohen’s between each pair of raters between 0.82-0.87. The error code distribution can be seen in Figure 2. The rubric is presented in Table 4.
| Rating | Definition of Rating |
| N/A | True/False does not apply here. Most stories, dialogues, or fictional works would be considered N/A, unless they are promoting a falsehood about the real world. |
| Not Sure | Claims are made that may not be true, but you aren’t sure. Choose this if it would take over 10 minutes to verify the factuality of the claim. |
| No Issues | All claims are factual and accurate. (Out-of-date is fine, e.g. “Barack Obama is the US President”) |
| Minor Issue(s) | There are small inaccuracies. E.g., it may be broadly correct but frame something in a misleading way. |
| Clear Issue(s) | There are clear mistakes in factuality. |
All ratings and rationales are made available. In addition, each datum in SmolDoc is given with a simple factuality annotation, which has the value has_errors if any one of the ratings was any of Minor Issues or Clear Issues, and ok otherwise. For some use-cases, like question-answering, practitioners may want to filter out nonfactual data; for others, like translation, one may not be troubled by factual errors. In addition to filtering, this also provides the first factuality dataset for most of these languages.
7 Conclusion
We have open-sourced the Smol dataset, a professionally-translated dataset covering 124 (and growing!) low resource languages and targeting the tasks of translation and factuality. It comprises SmolDoc and SmolSent, two training datasets with the complementary strengths of sentence selection (complex, and high token coverage) and document generation (contextual, varied domains, simpler sentences) respectively. We demonstrate that finetuning Gemini 2.0 Flash on these yields to substantial improvements in translation quality. Smol joins a growing body of resources to support underserved languages in the age of AI.
8 Limitations
The Smol data would benefit from a more thorough review, audit, and correction from community members outside of the translators who created it. Future work on Smol-like datasets should also focus on non-English source text that is not only maximally authentic in the given language, but also covers the topics and concepts most relevant to those languages. This approach is more difficult and would require significant work and review to do correctly. Finally, more research is needed to understand and prevent the overfitting that comes with multi-way parallel data.
Appendix A Operational Definition of LRL
In this paper, we operationally define LRL as any language beyond the first 100 supported by most traditional crawls and MT providers. Since there is some variation in which languages exactly this is, we concretize it as the set of 104 languages supported in Google Translate prior to 2020. These are the languages for which launchable quality was possible before LLM-type models like M4 (Arivazhagan et al., 2019; Bapna et al., 2022) and PaLM (Anil et al., 2023) came on these scene. It is also worth noting that since these languages were on the product for much longer, they have much more machine-translated content online from services that used Google Translate for internationalization. These 104 languages are: af am ar az be bg bn bs ca ceb co cs cy da de el en eo es et eu fa fi fil fr fy ga gd gl gu ha haw hi hmn hr ht hu hy id ig is it he ja jv ka kk km kn ko ku ky la lb lo lt lv mg mi mk ml mn mr ms mt my ne nl no ny pa pl ps pt ro ru sd si sk sl sm sn so sq sr st su sv sw ta te tg th tr uk ur uz vi xh yi yo zh zh-Hant zu.
Rightly speaking, the languages outside of this set might better be termed "Very Low-Resource" instead of just "Low Resource", since the 104 languages above do include languages like Hawaiian, Javanese, Yiddish, and Hmong, which are by no stretch of the imagination high-resource. We will leave more rigorous definitions to future work.
Appendix B SmolSent details
B.1 Evaluating the SmolSent selection process
In Section 3.1, we describe experiments used to validate the selection process for SmolSent. We train a backbone MT system is pretrained on the Madlad-400 dataset (Kudugunta et al., 2023). The following languages are held out of the training data to be used for fine-tuning experiments: Catalan, Icelandic, Marathi, Turkish, Maltese, Xhosa, Tamil, Basque, and Tajik. The model itself is a 1B parameter encoder-decoder Transformer and is trained from scratch on the Madlad-400 data. Each of the candidate data selection methods is used to select data from the held-out languages, and then each candidate set is used in turn to finetune the backbone model. The results of this finetune step are reported in Table 1, where the set-cover approach is shown to be most effective.
B.2 Notes on Researcher in the Loop
Researcher In the Loop extends the greedy set cover approach thusly: rather than always picking the highest-scoring sentence, we iteratively show the researcher a batch of the 20 highest scoring sentences according to several scores, and let the researcher pick and optionally edit each sentence at each iteration. At each iteration, the researcher may also remove any number of this batch’s sentences from the reservoir. Allowing the researcher to see and edit the sentences allows ensures that the sentences are of high-quality. To deal with the length bias issue, we showed not only sentences that maximize coverage percent, but also that maximize heuristics that weighted the coverage with the number of new tokens hit, like log(coverage_percent)*n_hits .
As described in the paper text, this approach is designed to combat issues such as honeypot sentences. Example Honeypot sentences can be seen in Table B.1
| Sentence |
| Individual determine can get prolonged, reduce along with attractive. |
| Sell hand mood situation connect proper decision today spread true. |
| Demand indeed off forget act special well treat sometimes notice. |
| Agree board book oh trust by attractive supply deal together. |
| Picture exactly could ability impact advance then same admire across. |
| One physically courage both information language issue laugh common. |
B.3 Corpus statistics on SmolSent
To measure “Bang for our Buck” we define the excess-token ratio as the number distinct tokens in the set cover divided by the number of target tokens, and use it along with the coverage percent to understand the SmolSent dataset. Table B.2 compares corpus statistics of SmolSent to four other corpora. sametoks picks a random set of sentences from CommonCrawl until it has the same number of tokens as SmolSent; this only covers 50% of the target tokens and has an excess-token-ratio ratio of 3.3, much worse than SmolSent’s value of 2.3. The samecov baseline randomly picks common-crawl sentences until it has the same token coverage as SmolSent, which necessitates a 67x larger set of sentences, and a correspondingly bloated excess token coverage ratio of 23.1. As a further reference we compare tiers 1 (largest) and 5 (smallest; comparative size to SmolSent) of SmolDoc. As expected of machine-generated text, they have a worse value, corresponding to a narrower spectrum of vocabulary used.
| set | N sent | toks | types | cov% | |
| SmolSent | 863 | 12k | 5.5k | 2.3 | 99.6% |
| sametoks | 863 | 12k | 3.8k | 3.3 | 50.4% |
| samecov | 57578 | 877k | 38k | 23.1 | 99.6% |
| SmolDoc-t1 | 6979 | 108k | 8.8k | 12.3 | 80.5% |
| SmolDoc-t5 | 820 | 12k | 2.8k | 4.3 | 40.2% |
Appendix C SmolDocDetails
C.1 SmolDoc data tier details
Per-tier statistics on the SmolDoc dataset can be seen in Table C.1.
| set | N langs | ex | tok | char | ex/LP | tok/LP | char/LP |
| SmolDoc-t1 | 5 | 2.9k | 537k | 3.0M | 584 | 107k | 604k |
| SmolDoc-t2 | 31 | 14k | 2.7M | 15M | 450 | 88k | 496k |
| SmolDoc-t3 | 24 | 6.7k | 1.2M | 6.8M | 280 | 50k | 281k |
| SmolDoc-t4 | 8 | 1.0k | 184k | 1.0M | 126 | 23k | 128k |
| SmolDoc-t5 | 38 | 2.5k | 448k | 2.5M | 66 | 12k | 65k |
C.2 Details on SmolDoc prompt creation
To avoid biases from overly tempted prompts, we put in significant effort to make sure the prompts were all very different. Each prompt drew at random from the following elements:
random selection of English words to use in the response
one of 600 manually created topics, e.g. “volcanic eruptions” or "A special tree”
one of 50 tone/tense categories, e.g. “Please use the subjunctive mood.”, “Use an effusive tone.”
A style prompt, e.g. “You are the author R.K. Narayan.” or “You are a mother talking to her son.”
A text modality, e.g. story/dialogue/essay
In addition to this, we added a few more sources of prompts:
Prompts based on urls, meant to simulate different web domains, like Wikipedia and reddit.
Prompts based on continuing the sentences from SmolSent
Prompts based on current events, history, and daily life in different countries
Special effort was made to include dialogues (to get more spoken register) and recipes (unique domain that may also be important to translate).
For each prompt, we generated 8 responses (T=0.7). These were ranked by their simple token density (unique tokens over total tokens), and the top two were chosen for consideration. Using the Researcher-in-the-loop mentality (“measure twice cut once!”), we went over 1000+ responses by hand and scored/edited them. This was mainly to filter out questionable or boring responses. A typical paragraph scored as 0 would be LLM-speak like “X is a complex and multifaceted problem with no easy solution. Here are some suggestions. Keep in mind that there is no one-size-fits all solution, and ultimately, the choice is up to you. […]”.
Example prompts can be seen in Table C.2.
| Example Prompts |
| You are Ernest Hemingway. Write a dialogue about road rage. Use a didactic tone. |
| Write a 1-paragraph story concerning an Irish wake. |
| You are a teenager talking to his friend. Please carefully craft a 1-paragraph bit about an engineer who subsists off coffee. Try to include the words “confirmed”, “move” and “above”. |
| Give a typical, yet interesting, example of something you would find on reddit. |
| Please write a few paragraphs about challenges facing Ethiopia. |
| Please write a long passage starting with ‘Mum and Dad pause their debate when we hear this creepy clacking that sounds like hail falling.’ |
| Write a recipe for baking an almond cake. |
C.3 SmolDoc Errata
Orthographies
Several languages use irregular orthographies. Most notable is Mooré/Mossi (mos), where different translators have used a variety of different conventions. After soliciting community feedback, we plan to release standardized versions to the data.
document selection
When collecting the data for SmolDoc for the Indian languages, we mistakenly included a variety of documents that fell below the corpus diversity threshold described in Section 3.2.
Appendix D N-shot: ChrF-counterweighted RAG
To have a strong baseline for N-shot results, we adopt a RAG-based approach that resembles the greedy set-cover algorithm. For each sentence in the eval set, we want the best coverage of the source sentence -grams as possible, with the least redundancy among exemplars. Therefore, we iteratively choose the exemplar whose source side has the minimum ChrF to the eval source. However, when counting the true positives in the ChrF calculation, we weight the count of each ngram by , where is the number of times has been seen among the exemplars chosen so far, and is a parameter to control how close this algorithm is to ngram set-cover. We use .The set of exemplars we choose from is the concatenation of SmolSent and SmolDocSplit.
Appendix E Prompts for Decoding
For 0-shot prompting, we used the following, fairly wordy prompt, the SL and TL standing for the source and target language name, respectively:
You are an expert translator. I am
going to give you some example
pairs of text snippets where the
first is in ${SL} and
the second is a translation of
the first snippet into
${TL}. The sentences
will be written
${SL}: <first sentence>
${TL}: <translated first sentence>
After the example pairs, I am
going to provide another sentence
in ${SL} and I want
you to translate it into
${TL}. Give only the
translation, and no extra
commentary, formatting, or
chattiness. Translate the text
from ${SL} to ${TL}.
For finetuned models, there is no need for such a wordy prompt, and indeed it only risks overfitting. Therefore, we used the following minimalist prompt:
Translate from ${SL} to ${TL}:
Appendix F Volunteer contributions
A few languages have extra details that need to be called out here.
F.1 Translations for Cantonese
A volunteer team of Cantonese speakers at Google pulled together to translate the maximal set of Smol text. Mingfei Lau and Jonathan Eng were the main leaders of this effort, and the contributors to translation and post-editing were (alphabetically): Tsz Yan Au, Emily Awesome, Jason Chan, Siu Man Chan, Vicky Chan, Yiwang Chen, Kinton Cheung, Mingo Choi, Andy Chow, Ashley Chow, Olivia Chow, Daniel (Ying Wai) Fan, Thomas Fung, Vikki Ha, Joshua Kwong, Liam Lee Pong Lam, Jonas Lau, Ying Tung (Grace) Law, Crystal Lee, Aki Leung, Derek Leung, Jackie Leung, Thomas Leung, Mu Li, Alicia Liu, Malena Loosli, Chui McConnell, Ken Ng, Nicholas Ng, Tonia Shen, Helen Shum, Franky Sze, Eric Tang, Tommy Tse, Daniel Wong, Danny Wong, Maggie Wong, Pinki Wong, Jeffrey Yu, Shanelle Yu, Shing Fung Yue, Miranda Zhang, and Willis Zhang.
F.2 Translations for NKo
The initial delivery for the NKo language (nqo) had a wide variety of errors. We reached out to the authors from Doumbouya et al. (2023), who did a complete re-translation of the text.
F.3 Translations for Zazaki, Hawrami, and Gilaki
Sina Ahmadi gratefully acknowledges support from the UZH Postdoc Grant (reference number 269093).
F.4 Translations for Zarma
Annotation Pipeline
The Zarma translation process of SMOL—all the subsets—was done through a combination of automatic and human in the loop methods. We leveraged some existing tools that our team developed to speed up the annotation process. We first used a baseline bidirectional model that we developed to produce initial translation of the samples. These machine translated samples were then passed through our Zarma grammatical error correction model. This model was built by pre-training gemma-2-9b on Zarma data and fine tuning the checkpoint on grammar error correction data set using Direct Preference Optimization (DPO) settings. The outputs from this stage—both languages side by side—were then given to our team of annotators for review.
The annotators were given some guidelines—in addition to the general guidelines from Smol—for the annotations. These guidelines include:
Word adaptation: rules for handling technical terms, proper nouns, and domain-specific vocabulary that might not have direct equivalents in Zarma. E.g: all the scientific/technical words remain unchanged; and words that have known french-ized equivalent in Zarma must be used in their french-ized forms (for better understandability).
Prioritize understandability: guidelines to prioritize understandability and fidelity over word-for-word translation. We instructed annotators to focus on creating translations that sound natural and widely understandable by Zarma speakers.
Language specific constraints: language specific guidelines that cannot be generalized.
The pipeline speeds up the process while maintaining the quality, since some of the outputs from the automatic stages were already correct.
Zarma Community Attitude Towards Tech
: The Zarma community—and the whole Niger in general—are very open minded regarding technology. When we started our very first resource creation for the Zarma language, we received positive feedback and even help from the community, as long as we developed an openly accessible solution for the community. For the SMOL annotation, that trust helps us to receive valuable help. For instance, a government based institution verbally promised to accompany any language preservation—machine learning focus in our case—if the outcome will be open-sourced for community usage.
F.5 Post-Edits for Mooré
Annotation Pipeline
The annotation process for Mooré did not involve any automated components; everything was annotated by humans. The annotation focuses more on the guidelines provided by Smol, in addition to some more as in the Zarma case.
Mooré Community Perspective
The Mooré community, similarly to the zarma community, are very open minded towards technology; especially if it touches cultural/language preservation. One main feedback we received from some elders (parents of one of the annotators) was a warning to ONLY USE standard Mooré orthography, not any equivalent. They want the language to be well documented according to the language standards.
F.6 Post-Edits for Indonesian
A volunteer team post-edited translations of smoldoc and smolsent datasets that had done by Gemini 2.5 Pro. The contributors to translation and post-editing were Muhammad Ravi Shulthan Habibi, David Anugraha, and Genta Indra Winata. The post-edits resulted in about 70% of the machine translations being changed.
Translators agreed that the system output was often too “formal", “stiff", or “awkward". The “formal" translations were furthermore not formal in an acceptible sense, but “too awkward and stiff, even for a more formal situations", as an annotator said. Each word choice might be correct and standard in Indonesian, but when combined in a sentence, the result sounded unnatural. Therefore, the majority of the post-edits focused on making the translations sound more natural.
Nonetheless, overall the system output was already quite reasonable in terms of register. In some cases, though, it leaned toward being too rigid. The post-edits tried to loosen that into a consistent “medium” range, but with some flexibility depending on the style of each sentence (sometimes slightly more formal, sometimes slightly less) so the overall text still feels natural and coherent.
F.7 Translations for Languages of the Russian Federation
Traditionally, speakers of hundreds of Cyrillic-based languages in the Russian Federation translate datasets via Russian. For the success of this project, I (Ali Kuzhuget) first funded a professional translation into Russian, engaging Andrey Anisimov as the main translator. The proofreading was conducted by Farhad Fatkullin, Vice-President of the National League of Translators, together with machine translation specialist David Dalé. I also oversaw formatting correctness and coordinated the overall translation workflow.
In parallel, I supervise the translation of the dataset from Russian into Tuvan, using a dedicated Telegram chatbot for large-scale dataset translation. This tool enables multiple rounds of validation and systematic assessment of translation quality. Currently, representatives of about a dozen Cyrillic languages are in the process of translating the Smol dataset into their own languages through Russian and/or English (for example, Tuvan, Bashkir, Chuvash, and others), ensuring both linguistic accuracy and cultural relevance.
Appendix G Full results
Full per-language results can be seen in Table LABEL:tab:full-results. Results are sorted by the , which is the ChrF of the Both model minus the ChrF of the finetuned Both model—in other words, how much the finetuning on Smol improved the baseline model.
Google Translate Languages and their cousins
As mentioned in the results section, some languages see only very small improvements from finetuning on Smol, and others even see losses. These are mainly either high-resource languages, or close relatives to higher-resource languages. In the full table LABEL:tab:full-results below, The superscript indicates a language supported by Google Translate at the time of these experiments; a superscript like means that this language is closely related to the Google-Translate-supported language xx. We only consider the 108 languages that were present on Google Translate at the time of this work.
| ee | Both | +36.1 | 3.0 | 37.7 | 37.6 | 39.1 | 39.2 | 37.8 | 40.0 | 39.5 | 7.5 | 42.7 | 40.7 |
| kr | Both | +10.8 | 17.3 | 25.6 | 25.9 | 28.1 | 28.8 | 22.7 | 26.3 | 20.3 | 22.2 | 32.6 | 31.0 |
| kg | Both | +9.2 | 34.9 | 46.9 | 36.8 | 44.1 | 43.2 | 43.2 | 47.0 | 37.8 | 29.1 | 50.2 | 3.4 |
| bem | Both | +7.3 | 40.0 | 44.8 | 44.7 | 47.3 | 49.2 | 43.3 | 47.7 | 42.3 | 33.3 | 49.7 | 41.8 |
| dyu | Both | +5.3 | 17.9 | 22.5 | 23.3 | 23.2 | 23.7 | 23.9 | 24.4 | 21.0 | 4.5 | 22.4 | 12.5 |
| din | Both | +4.6 | 20.3 | 23.8 | 22.9 | 24.9 | 25.1 | 23.3 | 25.9 | 21.4 | 1.6 | 25.1 | 26.5 |
| luo | Both | +4.1 | 37.4 | 39.1 | 41.1 | 41.5 | 42.0 | 39.1 | 42.0 | 39.6 | 36.1 | 41.3 | 39.5 |
| fon | Both | +3.6 | 21.3 | 24.3 | 23.9 | 24.9 | 25.3 | 20.4 | 23.7 | 23.8 | 1.9 | 25.9 | 24.2 |
| bm | Both | +3.4 | 30.8 | 28.6 | 35.2 | 34.2 | 34.1 | 34.0 | 36.2 | 33.9 | 9.0 | 35.7 | 32.2 |
| ak | Both | +2.6 | 35.5 | 36.1 | 37.8 | 38.1 | 38.2 | 34.4 | 38.1 | 37.3 | 32.2 | 34.5 | 33.3 |
| ln | Both | +2.5 | 46.8 | 48.1 | 48.4 | 49.3 | 49.3 | 44.6 | 48.3 | 46.5 | 45.2 | 46.4 | 45.7 |
| wo | Both | +1.1 | 30.3 | 30.0 | 30.4 | 31.4 | 31.6 | 31.4 | 32.2 | 30.7 | 29.8 | 36.2 | 30.9 |
| ff | Both | +0.9 | 25.0 | 24.4 | 26.4 | 25.9 | 26.5 | 25.7 | 26.1 | 25.2 | 2.5 | 25.9 | 27.1 |
| om | Both | -0.8 | 40.1 | 38.0 | 39.0 | 39.3 | 39.4 | 39.0 | 40.2 | 41.3 | 38.4 | 41.4 | 39.1 |
| lg | Both | -1.1 | 42.5 | 39.9 | 41.4 | 41.4 | 41.7 | 42.0 | 43.1 | 43.5 | 41.0 | 43.6 | 41.1 |
| ber | Both | -3.2 | 25.3 | 20.6 | 22.9 | 22.1 | 21.9 | 28.5 | 25.2 | 31.1 | 2.8 | 21.0 | 32.4 |
| trp | SmolDoc | +29.7 | 8.4 | 6.5 | 37.8 | 38.1 | 39.1 | 24.7 | 35.8 | 27.2 | 20.3 | 35.9 | - |
| mni-M. | SmolSent | +26.4 | 2.9 | 30.0 | 1.2 | 29.3 | 29.3 | 29.6 | 31.8 | 33.6 | 1.3 | 45.6 | 0.8 |
| gaa | Both | +23.1 | 22.7 | 44.5 | 44.4 | 45.8 | 47.4 | 34.7 | 44.0 | 40.9 | 6.6 | 48.3 | - |
| dov | Both | +21.1 | 19.1 | 39.2 | 38.3 | 40.2 | 40.6 | 19.2 | 39.5 | 18.2 | 8.7 | 41.7 | - |
| ahr | neither | +17.8 | 24.2 | 31.8 | 41.9 | 42.0 | 42.8 | 32.8 | 39.0 | 30.0 | 36.9 | - | - |
| sus | Both | +17.8 | 11.3 | 28.3 | 26.8 | 29.1 | 30.3 | 26.1 | 29.4 | 20.7 | 5.6 | 34.6 | - |
| nqo | Both | +17.5 | 0.2 | 17.9 | 17.1 | 17.7 | 17.5 | 17.1 | 17.9 | 17.2 | 1.1 | 19.1 | - |
| alz | Both | +15.5 | 16.9 | 31.5 | 30.5 | 32.4 | 33.4 | 25.3 | 30.6 | 26.9 | 8.9 | 36.6 | - |
| lu | Both | +13.8 | 27.6 | 37.5 | 39.3 | 41.4 | 42.2 | 27.2 | 37.0 | 34.8 | 21.9 | - | - |
| cgg | Both | +12.2 | 32.6 | 40.5 | 40.5 | 44.8 | 44.8 | 37.3 | 42.2 | 37.7 | 28.3 | 42.8 | - |
| ks-D. | neither | +11.7 | 14.9 | 26.6 | 18.8 | 26.6 | 27.7 | 23.8 | 27.1 | 21.5 | 19.2 | - | 21.1 |
| brx | SmolSent | +11.6 | 24.3 | 36.0 | 0.4 | 35.9 | 37.0 | 30.8 | 35.9 | 36.2 | 5.2 | - | - |
| mag | neither | +8.1 | 47.0 | 45.3 | 55.7 | 55.1 | 54.7 | 47.3 | 51.8 | 47.4 | 48.7 | - | 59.4 |
| ki | Both | +7.7 | 32.6 | 38.2 | 38.0 | 40.3 | 40.6 | 35.9 | 42.0 | 39.1 | 10.5 | - | 38.4 |
| aa | Both | +7.4 | 14.2 | 20.1 | 20.3 | 21.6 | 21.8 | 18.9 | 20.6 | 18.9 | 5.6 | 23.1 | - |
| ks | neither | +7.3 | 22.1 | 29.4 | 0.4 | 29.4 | 29.7 | 28.0 | 30.4 | 30.5 | 26.3 | - | 36.7 |
| nr | neither | +7.0 | 48.9 | 54.0 | 51.1 | 55.9 | 57.5 | 48.0 | 53.6 | 51.2 | 45.5 | 59.5 | - |
| doi | neither | +6.6 | 34.3 | 28.1 | 41.4 | 40.9 | 41.3 | 35.9 | 39.5 | 38.2 | 27.7 | 40.4 | - |
| sat-L. | SmolSent | +6.4 | 12.8 | 19.5 | 15.5 | 19.2 | 20.8 | 25.3 | 22.5 | 22.7 | 21.3 | 22.7 | - |
| mfe | neither | +5.3 | 59.5 | 65.4 | 62.6 | 64.8 | 66.9 | 59.6 | 65.0 | 59.8 | 59.5 | 67.5 | - |
| ach | Both | +5.2 | 33.2 | 43.1 | 32.5 | 38.4 | 39.2 | 32.4 | 37.3 | 35.1 | 23.8 | 43.2 | - |
| ayl | neither | +4.5 | 47.3 | 51.7 | 51.9 | 51.8 | 53.9 | 45.3 | 48.9 | 46.3 | 48.6 | - | - |
| st | neither | +4.5 | 49.9 | 54.0 | 55.2 | 54.4 | 55.0 | 49.4 | 57.0 | 53.1 | 49.2 | 49.0 | 47.2 |
| ber-L. | Both | +4.2 | 26.1 | 27.9 | 30.3 | 30.3 | 30.9 | 27.6 | 32.7 | 32.1 | 21.2 | 34.7 | - |
| apd-S. | neither | +3.6 | 42.3 | 50.2 | 43.3 | 45.9 | 47.1 | 43.2 | 45.0 | 42.5 | 45.6 | - | - |
| ve | neither | +3.5 | 47.9 | 50.0 | 48.7 | 51.4 | 52.2 | 50.2 | 53.1 | 52.7 | 43.9 | 56.8 | - |
| kri | neither | +2.8 | 31.5 | 34.2 | 31.8 | 34.3 | 34.7 | 34.5 | 33.5 | 30.7 | 34.9 | 34.9 | - |
| tiv | Both | +2.5 | 23.8 | 25.7 | 25.8 | 26.3 | 26.5 | 22.3 | 24.2 | 24.5 | 1.5 | 25.2 | - |
| gn | SmolSent | +2.3 | 37.4 | 37.8 | 30.4 | 39.7 | 38.4 | 36.0 | 38.0 | 36.4 | 35.6 | 38.4 | 38.5 |
| mos | Both | +2.3 | 18.2 | 20.9 | 18.9 | 20.5 | 21.1 | 24.3 | 25.0 | 20.9 | 1.3 | - | 23.8 |
| tum | neither | +1.1 | 40.8 | 39.5 | 42.4 | 41.9 | 42.8 | 40.0 | 42.7 | 43.8 | 37.7 | 45.4 | 36.2 |
| ti | neither | +0.8 | 24.2 | 24.3 | 24.9 | 25.0 | 25.7 | 25.0 | 26.2 | 26.1 | 9.3 | 26.1 | 25.5 |
| yo | neither | +0.6 | 34.6 | 33.8 | 35.4 | 35.2 | 35.8 | 29.2 | 36.8 | 26.7 | 27.6 | 21.3 | 32.6 |
| tn | neither | +0.2 | 52.5 | 50.8 | 51.9 | 52.7 | 53.2 | 50.1 | 51.7 | 53.3 | 36.8 | 55.6 | 53.0 |
| ar-M. | neither | +0.1 | 40.1 | 41.8 | 39.1 | 40.2 | 40.9 | 40.5 | 40.8 | 40.4 | 41.0 | - | 43.0 |
| am | neither | +0.1 | 34.0 | 33.2 | 33.0 | 34.1 | 33.5 | 31.6 | 32.6 | 35.8 | 29.6 | 34.7 | 30.3 |
| ig | neither | -0.1 | 47.2 | 47.1 | 47.0 | 47.1 | 47.8 | 43.9 | 46.2 | 47.8 | 46.2 | 47.6 | 46.6 |
| so | neither | -0.1 | 49.7 | 46.2 | 50.0 | 49.6 | 49.1 | 48.7 | 49.8 | 50.3 | 50.8 | 50.6 | 48.6 |
| arz | neither | -0.1 | 48.6 | 46.1 | 48.9 | 48.5 | 47.8 | 49.6 | 50.3 | 48.8 | 49.7 | - | 49.6 |
| kl | SmolSent | -0.3 | 40.6 | 39.7 | 30.2 | 40.3 | 41.2 | 42.2 | 43.1 | 41.2 | 41.6 | 42.9 | - |
| sa | SmolSent | -0.3 | 33.0 | 32.9 | 26.7 | 32.7 | 33.4 | 31.8 | 33.0 | 32.1 | 32.0 | 35.2 | 29.0 |
| ay | SmolSent | -0.5 | 32.7 | 31.8 | 24.2 | 32.2 | 32.4 | 33.4 | 33.2 | 32.9 | 30.0 | 34.7 | 31.7 |
| sn | neither | -0.5 | 50.5 | 48.3 | 50.2 | 50.0 | 50.3 | 46.8 | 48.8 | 51.8 | 50.3 | 49.2 | 48.2 |
| efi | Both | -0.6 | 14.7 | 14.5 | 14.3 | 14.1 | 14.2 | 15.3 | 15.1 | 15.0 | 2.2 | - | - |
| ss | neither | -0.6 | 50.2 | 48.8 | 48.2 | 49.6 | 50.3 | 49.6 | 51.2 | 51.8 | 46.0 | 56.3 | 48.1 |
| yue | neither | -0.7 | 26.8 | 25.8 | 25.1 | 26.1 | 26.1 | 28.2 | 28.2 | 27.5 | 31.6 | 25.9 | 22.6 |
| bci | Both | -0.7 | 23.2 | 22.2 | 21.8 | 22.5 | 22.9 | 17.1 | 20.7 | 27.6 | 1.0 | 29.8 | - |
| ndc-Z. | neither | -1.0 | 29.2 | 27.9 | 28.9 | 28.2 | 28.3 | 27.6 | 28.0 | 29.6 | 28.6 | 29.5 | - |
| es | neither | -1.1 | 62.4 | 61.3 | 51.6 | 61.3 | 61.3 | - | - | 63.0 | - | 63.5 | 61.8 |
| sat | SmolSent | -1.3 | 32.4 | 30.9 | 1.0 | 31.1 | 30.8 | 34.7 | 36.0 | 36.3 | 1.8 | 35.7 | - |
| rw | neither | -1.4 | 45.2 | 43.1 | 43.0 | 43.8 | 43.8 | 43.2 | 44.0 | 45.1 | 44.7 | 48.8 | 43.4 |
| nd | neither | -1.4 | 43.9 | 41.5 | 42.6 | 42.5 | 43.2 | 42.3 | 42.9 | 44.5 | 43.6 | - | - |
| sw | neither | -1.5 | 66.7 | 64.6 | 64.2 | 65.2 | 64.8 | 64.0 | 65.5 | 67.2 | 66.5 | 65.3 | 60.5 |
| mg | neither | -1.9 | 52.8 | 48.9 | 51.6 | 50.9 | 51.5 | 52.4 | 52.5 | 53.3 | 52.2 | 52.6 | 52.1 |
| qu | SmolSent | -1.9 | 34.7 | 32.8 | 30.4 | 32.8 | 33.0 | 35.3 | 35.1 | 34.0 | 22.0 | 36.3 | 27.9 |
| zu | neither | -2.0 | 58.3 | 56.4 | 55.3 | 56.3 | 56.1 | 54.1 | 55.5 | 58.5 | 57.5 | 57.6 | 57.6 |
| lus | SmolSent | -2.1 | 42.6 | 39.7 | 38.0 | 40.5 | 41.4 | 40.6 | 41.5 | 43.8 | 33.5 | 42.6 | 39.0 |
| scn | neither | -2.2 | 52.4 | 47.3 | 50.0 | 50.2 | 50.8 | 49.9 | 51.4 | 52.1 | 49.5 | 53.3 | 51.0 |
| nso | neither | -2.3 | 46.8 | 42.8 | 43.6 | 44.5 | 44.6 | 46.9 | 47.7 | 48.1 | 46.9 | 47.6 | 45.5 |
| xh | neither | -2.3 | 53.9 | 49.7 | 50.7 | 51.6 | 51.8 | 51.6 | 52.3 | 53.9 | 53.7 | 54.8 | 51.2 |
| ne | neither | -2.7 | 54.3 | 51.8 | 51.7 | 51.6 | 52.1 | 52.4 | 52.5 | 52.4 | 52.7 | 54.9 | 45.2 |
| pa-A. | neither | -3.0 | 38.1 | 35.8 | 0.3 | 35.1 | 35.7 | 41.6 | 41.2 | 36.7 | 37.3 | 43.5 | - |
| aeb | neither | -3.3 | 46.5 | 41.9 | 42.3 | 43.2 | 43.4 | 45.9 | 46.8 | 47.6 | 49.2 | - | 43.8 |
| ha | neither | -3.4 | 54.5 | 50.7 | 49.9 | 51.1 | 51.5 | 50.9 | 51.0 | 54.1 | 53.9 | 53.8 | 53.9 |
| ts | neither | -3.6 | 50.6 | 47.2 | 46.4 | 47.0 | 48.1 | 49.7 | 50.1 | 51.6 | 49.0 | 52.9 | 51.3 |
| rn | neither | -3.9 | 44.5 | 39.3 | 40.5 | 40.6 | 40.9 | 43.1 | 43.4 | 46.2 | 44.3 | 45.4 | 45.0 |
| af | neither | -4.2 | 71.9 | 68.8 | 67.9 | 67.7 | 68.3 | 71.7 | 72.5 | 72.1 | 71.8 | 71.5 | 68.6 |
| bo | SmolSent | -4.3 | 41.3 | 36.7 | 34.7 | 37.0 | 37.3 | 42.6 | 42.1 | 43.3 | 19.8 | 41.8 | 36.9 |
| ny | neither | -5.1 | 55.0 | 47.5 | 50.5 | 49.9 | 49.7 | 53.0 | 53.1 | 55.3 | 53.9 | 55.8 | 50.3 |
| pcm | neither | -6.5 | 47.9 | 43.5 | 39.4 | 41.4 | 41.6 | 51.3 | 45.7 | 49.8 | 56.0 | - | - |
| tcy | SmolDoc | -6.8 | 34.7 | 22.0 | 28.1 | 27.9 | 28.8 | 28.2 | 29.3 | 36.7 | 21.6 | 39.1 | - |
| ktu | Both | -9.4 | 56.6 | 59.3 | 40.4 | 47.2 | 51.3 | 45.8 | 48.4 | 57.8 | 22.3 | 64.3 | - |